### v1

# Segment Anything Meets Point Tracking

Abstract
The Segment Anything Model (SAM) has established itself as a powerful zero-shot image segmentation model, employing interactive prompts such as points to generate masks. This paper presents SAM-PT, a method extending SAM’s capability to tracking and segmenting anything in dynamic videos. SAM-PT leverages robust and sparse point selection and propagation techniques for mask generation, demonstrating that a SAM-based segmentation tracker can yield strong zero-shot performance across popular video object segmentation benchmarks, including DAVIS, YouTube-VOS, and MOSE. Compared to traditional object-centric mask propagation strategies, we uniquely use point propagation to exploit local structure information that is agnostic to object semantics. We highlight the merits of point-based tracking through direct evaluation on the zero-shot open-world Unidentified Video Objects (UVO) benchmark. To further enhance our approach, we utilize K-Medoids clustering for point initialization and track both positive and negative points to clearly distinguish the target object. We also employ multiple mask decoding passes for mask refinement and devise a point re-initialization strategy to improve tracking accuracy. Our code integrates different point trackers and video segmentation benchmarks and will be released at https://github.com/SysCV/sam-pt.
{strip}
Figure 1: Segment Anything Meets Point Tracking (SAM-PT). SAM-PT is the first method to utilize sparse point propagation for Video Object Segmentation (VOS). The essence of SAM-PT is to extend SAM [16] with long-term point trackers to effectively operate on videos in a zero-shot manner. SAM-PT takes a video as input together with annotations of the target object in the first frame. These annotations are called “query points” and denote either the target object (positive points) or designate non-target segments (negative points). The points are tracked throughout the video using point trackers that propagate the query points to all video frames, producing predicted trajectories and occlusion scores. SAM is subsequently prompted with the non-occluded points in the trajectories as to output a segmentation mask for each video frame independently.
1 Introduction
Video segmentation benefits a myriad of applications, including autonomous driving, robotics, and video editing. Despite significant progress made in the past few years with deep neural networks [41, 4, 37, 3], the current methodologies falter when faced with unseen data, particularly in zero-shot settings. These models struggle to maintain consistent performance across diverse scenarios without specific video segmentation data for fine-tuning.
The prevailing methods [4, 3] in semi-supervised Video Object Segmentation (VOS) and Video Instance Segmentation (VIS) exhibit performance gaps when dealing with unseen data, particularly in a zero-shot setting, i.e., when these models are transferred to video domains they have not been trained on and encompass object categories that fall outside of the training distribution.
A potential route towards overcoming these challenges lies in adapting successful models in the image segmentation domain for video segmentation tasks. One such promising model is the Segment Anything Model (SAM) [16]. SAM is a powerful foundation model for image segmentation, trained on the large-scale SA-1B dataset, which contains an astounding 11 million images and over 1 billion masks. This extensive training set enables SAM’s impressive zero-shot generalization capabilities. The model is highly adaptable, able to produce high-quality masks from single foreground points, and has demonstrated robust performance across a range of downstream tasks under zero-shot transfer protocols. While SAM demonstrates powerful zero-shot capabilities for image segmentation, it is not innately suited for video segmentation tasks.
Recent efforts have been made to adapt SAM for video segmentation. For instance, TAM [40] integrates SAM with the state-of-the-art memory-based mask tracker XMem [4]. Likewise, SAM-Track [6] combines SAM with DeAOT [41]. While these methods mostly recover the performance on in-distribution data, they fall short in preserving the original performance of SAM in more challenging, zero-shot settings. Other methods that do not leverage SAM, such as SegGPT [36], can successfully solve a number of segmentation problems using visual prompting, but still require mask annotation for the first video frame. This problem represents a significant barrier in zero-shot video segmentation, particularly as we seek to develop methods that can easily generalize to unseen scenarios and consistently deliver high-quality segmentation across diverse video domains.
We introduce SAM-PT (Segment Anything Meets Point Tracking), depicted in Segment Anything Meets Point Tracking. This is the first method to utilize sparse point tracking combined with SAM for video segmentation, offering a new perspective on solving the problem. Instead of employing object-centric dense feature matching or mask propagation, we propose a point-driven approach that capitalizes on tracking points using rich local structure information embedded in videos. As a result, it only requires sparse points annotation to denote target object in the first frame and provides better generalization to unseen objects, a strength demonstrated on the open-world UVO [34] benchmark. This approach also helps preserve the inherent flexibility of SAM while extending its capabilities effectively to video segmentation.
SAM-PT prompts SAM with sparse point trajectories predicted using state-of-the-art point trackers, such as PIPS [11], harnessing their versatility for video segmentation. We identify that initializing points to track using K-Medoids cluster centers from a mask label was the strategy most compatible with prompting SAM. Tracking both positive and negative points enables the clear delineation of target objects from their background. To further refine the output masks, we propose multiple mask decoding passes that integrate both types of points. In addition, we devised a point re-initialization strategy that increases tracking accuracy over time. This approach involves discarding points that have become unreliable or occluded, and adding points from object parts or segments that become visible in later frames, such as when the object rotates.
Notably, our experimental results highlight that SAM-PT competes with existing zero-shot methods [36] or outperforms them [45, 2, 12, 35] on several video segmentation benchmarks. This comes without the need for any video segmentation data during training, underscoring the robustness and adaptability of our approach. SAM-PT holds the potential to enhance progress in video segmentation tasks, particularly in zero-shot scenarios.
2 Related Work
Point Tracking for Video Segmentation.
Classical feature extraction and tracking methods such as Lucas-Kanade [23], Tomasi-Kanade [31], Shi-Tomasi [29], SIFT [22], and SURF [1], as well as newer methods such as LIFT [43], SuperPoint [7], and SuperGlue [28], have all demonstrated proficiency in identifying or tracking sparse features and establishing long-range correspondences. However, their effectiveness is confined to a specific set of distinct interest points and they often struggle when applied to non-rigid, dynamic scenes. Flow-based methods, such as RAFT [30], excel in tracking dense points between successive frames. However, they stumble with deriving accurate long-range point trajectories. When chaining flow predictions over time, errors tend to accumulate and lead to drift, while occlusions result in tracking failures.
Significant strides have recently been made in long-term point tracking across video frames, as evinced by methods such as TapNet [9] and PIPS [11], as well as the concurrent and state-of-the-art OmniMotion [32] and TAPIR [10] techniques. These approaches optimize long-range point trajectories across an entire video, navigating mostly well through periods of occlusion.
Our work stands apart as the first to integrate these successful long-term point tracking methods, utilizing them to guide a promptable foundation model for image segmentation toward performing video segmentation tasks.
Segment and Track Anything models.
SAM [16] is an innovative image segmentation model for promptable image segmentation, trained on over 1 billion segmentation masks. It showcases remarkable zero-shot generalization abilities and can produce high-quality masks from a single foreground point. To further improve the quality of the masks, especially when segmenting objects with intricate structures, HQ-SAM [15] extends SAM with a learnable high-quality output token which proves efficient in diverse segmentation domains. However, SAM and HQ-SAM cannot be directly used to solve video segmentation tasks.
A few concurrent works extend SAM, for example, TAM [40] and SAM-Track [6] combine SAM with state-of-the-art mask trackers (such as XMem [4] and DeAOT [42]) to perform interactive video object segmentation. These methods employ SAM for mask initialization or correction and XMem/DeAOT for mask tracking and prediction. Using the pre-trained mask trackers recovers the in-distribution performance, but hinders the performance in zero-shot settings. PerSAM [45] also demonstrates the ability to track multiple reference objects in a video. Instead of building an interactive tracking pipeline or SAM fine-tuning, we focus on learning robust associations for diverse objects in zero-shot scenarios.
Zero-shot VOS / VIS.
Among the non-SAM-based methods, Painter [35] and its SegGPT [36] extension are another sort of generalist models for solving a variety of image and segmentation tasks. These methods likewise use visual prompting techniques but are inherently different frameworks from SAM. Despite its wide applicability, Painter shows lacking performance in video segmentation tasks. Conversely, SegGPT successfully uses in-context prompting to achieve one-shot video object segmentation performance comparable to ours, also without training on any video data. The training domains, however, notably differ between SegGPT and our method.
STC [12] and DINO [2] also do not use any video segmentation data during training. In the semi-supervised video object segmentation, they take a reference mask as input and perform frame-by-frame feature matching, which propagates the reference mask across the entirety of the video. Our SAM-PT, on the other hand, diverges substantially from these methodologies by adopting point tracking, eschewing the process of frame-by-frame feature matching. Additionally, our method requires only sparse points to represent the target object, rather than a full reference mask, and yields superior performance on conventional semi-supervised video object segmentation benchmarks.
3 Method
We propose SAM-PT to adapt SAM, a foundation model for image segmentation, for addressing video segmentation tasks in a zero-shot setting. SAM-PT combines the strengths of existing prominent point trackers, such as PIPS [11] and TapNet [9], with the powerful image segmentation of SAM to enable tracking of anything in videos. First, Sec. 3.1 briefly describes SAM. Sec. 3.2 then introduces our SAM-PT method with its four constituent steps. Finally, Sec. 3.3 analyzes and highlights the method’s novelty as the first point-driven video segmentation method compared to existing works.
3.1 Preliminaries: SAM
The Segment Anything Model (SAM) [16] is a novel vision foundation model designed for promptable image segmentation. SAM is trained on the large-scale SA-1B dataset, which contains 11 million images and over 1 billion masks. SA-1B has 400 times more masks than any existing segmentation dataset. This extensive training set facilitates SAM’s impressive zero-shot generalization capabilities to new data. SAM has showcased its ability to produce high-quality masks from a single foreground point, and has demonstrated robust generalization capacity on a variety of downstream tasks under a zero-shot transfer protocol using prompt engineering. These tasks include, but are not limited to, edge detection, object proposal generation, and instance segmentation.
SAM comprises of three main components: an image encoder, a flexible prompt encoder, and a fast mask decoder. The image encoder is a Vision Transformer (ViT) backbone and processes high-resolution $1024\times 1024$ images to generate an image embedding of $64\times 64$ spatial size. The prompt encoder takes sparse prompts as input, including points, boxes, and text, or dense prompts such as masks, and translates these prompts into $c$-dimensional tokens. The lightweight mask decoder then integrates the image and prompt embeddings to predict segmentation masks in real-time, allowing SAM to adapt to diverse prompts with minimal computational overhead.
3.2 Ours: SAM-PT
While SAM shows impressive capabilities in image segmentation, it is inherently limited in handling video segmentation tasks. Our Segment Anything Meets Point Tracking (SAM-PT) approach effectively extends SAM to videos, offering robust video segmentation without requiring training on any video segmentation data.
SAM-PT is illustrated in Fig. 2 and is primarily composed of four steps: 1) selecting query points for the first frame; 2) propagating these points to all video frames using point trackers; 3) using SAM to generate per-frame segmentation masks based on the propagated points; 4) optionally reinitializing the process by sampling query points from the predicted masks. We next elaborate on these four steps.
Figure 2: Segment Anything Meets Point Tracking (SAM-PT) overview. The essence of SAM-PT is to extend image segmentation foundation model to effectively operate on videos. SAM-PT has four steps: $\mathbf{1}$) Query Points Selection. It starts with first-frame query points which denote the target object (positive points) or designate non-target segments (negative points). These points are provided by the user or derived from a ground truth mask. $\mathbf{2}$) Point Tracking. Initiated with the query points, our approach leverages point trackers to propagate the points across video frames, predicting point trajectories and occlusion scores. $\mathbf{3}$) Segmentation. The trajectories are then used to prompt the Segment Anything Model (SAM) and output per-frame mask predictions. $\mathbf{4}$) Point Tracking Reinitialization. Optionally, the predicted masks are used to reinitialize the query points and restart the process when reaching a prediction horizon $h$. Re-initialization helps by getting rid of unreliable and occluded points, and adds points from object parts or segments that become visible in later frames, such as when the object rotates.
1) Query Points Selection.
The process begins with defining query points in the first video frame, which either denote the target object (positive points) or designate the background and non-target objects (negative points). Users can manually and interactively provide query points, or they may be derived from a ground truth mask. For example, in the case of semi-supervised video object segmentation, the ground truth mask is provided for the first frame where the object appears. We derive the query points from ground truth masks using different point sampling techniques by considering their geometrical locations or feature dissimilarities, as depicted in Fig. 3. These sampling techniques are:
  • Random Sampling: An intuitive approach where query points are randomly selected from the ground truth mask.
  • K-Medoids Sampling: This technique takes the cluster centers of K-Medoids clustering [26] as query points to ensure good coverage of different parts of the object and robustness to noise and outliers.
  • Shi-Tomasi Sampling: This method extracts Shi-Tomasi corner points from the image under the mask as they have been shown to be good features to track [29].
  • Mixed Sampling: A hybrid method combining the above techniques since it might benefit from the unique strengths of each.
While each method contributes distinct characteristics that influence the model’s performance, our ablation study reveals that K-Medoids sampling yields the best results with a good full coverage of various segments of the complete object. Shi-Tomasi sampling follows closely, indicating their respective strengths in this context. The selection and arrangement of these points considerably affect the overall video segmentation performance, thus determining the optimal method is crucial.
Figure 3: Positive Point Sampling. For an image paired with either a ground truth or predicted segmentation mask, positive points are sampled from within the mask area using one of the following point sampling methods: Random, K-Medoids [26], Shi-Tomasi [29], or Mixed. Notably, Random Sampling and K-Medoids Sampling only require the segmentation mask for input, not the corresponding input image. For negative points, we always use Mixed Sampling on the target object’s background mask.
2) Point Tracking.
Initiated with the query points, we employ robust point trackers to propagate the points across all frames in the video, resulting in point trajectories and occlusion scores. We adopt the state-of-the-art point tracker PIPS [11] to propagate the points as PIPS shows moderate robustness toward long-term tracking challenges such as object occlusion and re-appearance. This is also shown more effective than methods such as chained optical flow propagation or first-frame correspondences in our experiment section.
3) Segmentation.
In the predicted trajectories, the non-occluded points serve as indicators of where the target object is throughout the video. This allows us to use the non-occluded points to prompt SAM, as illustrated in Fig. 4, and leverage its inherent generalization ability to output per-frame segmentation mask predictions. Unlike conventional tracking methods that require training or fine-tuning on video segmentation data, our approach excels in zero-shot video segmentation tasks.
We combine positive and negative points by calling SAM in two passes. In the initial pass, we prompt SAM exclusively with positive points to define the object’s initial localization. Subsequently, in the second pass, we prompt SAM with both positive and negative points along with the previous mask prediction. Negative points provide a more nuanced distinction between the object and the background and help by removing wrongly segmented areas.
Lastly, we execute a variable number of mask refinement iterations by repeating the second pass. This utilizes SAM’s capacity to refine vague masks into more precise ones. Based on our ablation study, this step notably improves video object segmentation performance.
Figure 4: Interacting with SAM in SAM-PT. In the first pass, SAM is prompted exclusively with positive points to define the object’s initial localization. In the second pass, both positive and negative points along with the previous mask prediction are fed to the same mask decoder for further mask refinement. The negative points remove segments from the background and neighboring objects and notably help in cases when the point tracker mistakenly predicts positive points off the target object. The second pass is repeated iteratively to get a refined segmentation mask.
4) Point Tracking Reinitialization.
We optionally execute a reinitialization of the query points using the predicted masks once a prediction horizon of $h=8$ frames is reached, and denote the variant as SAM-PT-reinit. Upon reaching this horizon, we have $h$ predicted masks and will take the last predicted mask to sample new points. At this stage, all previous points are discarded and substituted with the newly sampled points. Following this, steps 1) through 4) are repeated with the new points, starting from the horizon timestep where reinitialization occurs. The steps are iteratively executed until the entire video is processed. The reinitialization process serves to enhance tracking accuracy over time by discarding points that have become unreliable or occluded, while incorporating points from object segments that become visible later in the video. Other reinitialization variants are discussed in Appendix A and included in the ablation study in Sec. 4.3.
3.3 SAM-PT vs. Object-centric Mask Propagation
With sparse point tracking combined with prompting SAM, SAM-PT distinguishes itself from traditional video segmentation methods that depend on dense object mask propagation, as noted in Tab. 1. To propagate the first-frame GT label to the remaining video frames, traditional techniques commonly use feature matching with masks cached to a mask memory [40, 6, 4, 41], frame-by-frame feature matching [12, 2], feature matching with the first-frame mask [45], optical flow [39], and, recently, in-context visual prompting [35, 36]. In contrast, SAM-PT introduces a unique approach to video object segmentation, employing the robust combination of point tracking with SAM, which is inherently designed to operate on sparse point prompts.
The point propagation strategy of SAM-PT offers several advantages over traditional object-centric tracking methods. First, point propagation exploits local structure context that is agnostic to global object semantics. This enhances our model’s capability for zero-shot generalization, an advantage that, coupled with SAM’s inherent generalization power, allows for tracking diverse objects in diverse environments, such as on the UVO benchmark. Second, SAM-PT allows for a more compact object representation with sparse points, capturing enough information to characterize the object’s segments/parts effectively. Finally, the use of points is naturally compatible with SAM, an image segmentation foundation model trained to operate on sparse point prompts, offering an integrated solution that aligns well with the intrinsic capacities of the underlying model.
Comparing SAM-PT with conventional methods in Tab. 1, SAM-PT emerges as superior or comparable to methods that refrain from utilizing video segmentation data during training. However, there is a performance gap that exists between such methods and those that leverage video segmentation training data in the same domain, such as XMem [4] or DeAOT [41]. Further, the potential of our model extends beyond video object segmentation to other tasks, such as Video Instance Segmentation (VIS), thanks to the inherent flexibility of our point propagation strategy.
In summary, SAM-PT is the first method that introduces sparse point propagation combined with prompting a image segmentation foundation model to perform zero-shot video object segmentation. It provides a fresh perspective and adds a new dimension to the study of video object segmentation.
Table 1: Comparison of semi-supervised Video Object Segmentation (VOS) methods with respect to mask annotation requirements and propagation techniques. Our method, SAM-PT, is the first method for VOS that uses sparse point propagation. With such compact mask representation, we achieve the highest $\mathcal{J\&F}$ scores on the DAVIS $2016$ and $2017$ validation subsets among methods that do not utilize any video segmentation data during training. The methods are compared based on their use of video mask data during training, whether they are evaluated zero-shot on DAVIS, what first-frame labels they require, and what label propagation technique they employ.
Method Video Mask Zero- Shot Frame Init. Propagation DAVIS $\mathbf{2016}$ DAVIS $\mathbf{2017}$
SiamMask [33] Box Feature Correlation $69.8$ $56.4$
QMRA [19] Box Feature Correlation $85.9$ $71.9$
TAM [40] Points Feature Matching $88.4$
SAM-Track [6] Points Feature Matching $92.0$
XMem [4] Mask Feature Matching $92.0$ 87.7
DeAOT [41] Mask Feature Matching $\mathbf{92.9}$ $86.2$
Painter [35] Mask Mask Prompting - $34.6$
STC [12] Mask Feature Matching - $67.6$
DINO [2] Mask Feature Matching - $71.4$
PerSAM-F [45] Mask Feature Matching - $71.9$
SegGPT [36] Mask Mask Prompting $82.3$ $75.6$
SAM-PT (ours) Points Points Prompting $\mathbf{83.1}$ $\mathbf{76.6}$
4 Experiments
4.1 Datasets
In the following subsections, we present an overview of the datasets used in our study. Section 4.1.1 provides a brief introduction to the Video Object Segmentation task and outlines the specific datasets we utilize for this task. Similarly, Section 4.1.2 discusses the Video Instance Segmentation task and the dataset associated with it.
4.1.1 Video Object Segmentation
Video Object Segmentation (VOS) refers to the process of segmenting a specific object across an entire video sequence. Semi-supervised VOS (also known as one-shot VOS or semi-automatic VOS) is the primary setting for VOS on which we evaluate our method. In this setting, the ground truth object mask of the first frame is provided, and the task is to predict the masks for subsequent frames. Alternatively, the first frame label can be a bounding box instead of a segmentation mask, or a set of points as is the case for our method. We evaluate our method on four VOS datasets: DAVIS $2016$, DAVIS $2017$ [27], YouTube-VOS $2018$ [38], and MOSE $2023$ [8].
DAVIS $\mathbf{2016}$ [27].
DAVIS $2016$ is a single-object VOS benchmark, consisting of $20$ highly diverse video sequences, each of which possesses well-annotated segmentation masks.
DAVIS $\mathbf{2017}$ [27].
A multi-object extension of its $2016$ version, DAVIS $2017$ includes $60$ videos in the training set and $30$ videos in the validation set, comprising a total of $197$ different objects. The video scenarios within this dataset are small but diverse.
YouTube-VOS $\mathbf{2018}$ [38].
YouTube-VOS $2018$ is a large-scale dataset collected from YouTube, comprising $3471$ training videos encompassing $65$ categories and $474$ validation videos with an additional $26$ unseen categories. The diversity in categories and the inclusion of seen and unseen classes allow for a comprehensive evaluation of a given model’s generalization capability.
MOSE $\mathbf{2023}$ [8].
MOSE $2023$ is a recently introduced dataset designed for multiple object segmentation and tracking in complex scenes. This dataset is replete with challenges such as the transient visibility of objects, the presence of minute or less noticeable entities, extensive occlusions, and scenes with a high object density. By design, each video in this dataset must contain multiple objects so that occlusions must be present, and objects must show sufficient motion, as opposed to being stationary or showing little movement.
Metrics.
We report the standard evaluation metrics for video object segmentation [8, 27, 38], including region similarity $\mathcal{J}$, contour accuracy $\mathcal{F}$, and their average, $\mathcal{J\&F}$.
4.1.2 Video Instance Segmentation
Video Instance Segmentation (VIS) is a task that combines object detection, instance segmentation, and object tracking across video frames, which aims to identify and segment each object instance over the whole video sequence. This is a much less explored task compared to VOS but has been gaining interest. We evaluate our method on the dense-video task of the UVO v1.01.01.01.0 [34] dataset.
[…]
The Unidentified Video Objects (UVO) dataset is designed to recognize and segment all objects regardless of the categories, even those unseen during training, thereby focusing on VIS in the open world. Each video in UVO features on average 12.3 object annotations, a considerable increase from previous datasets having only 2 or 3 objects per video on average. UVO sources its videos from the Kinetics-400 [14] dataset and contains three different splits: FrameSet, VideoSparseSet, and VideoDenseSet. The VideoDenseSet consists of 3-second clips annotated densely at 30fps and tracked over time. The primary goal of VideoDenseSet is to study video open-world segmentation. Objects identifiable under COCO categories carry their respective COCO labels, while ambiguous objects or those outside the COCO taxonomy are labeled as “other”. This meticulous and exhaustive annotation structure makes the VideoDenseSet ideal for research areas that require an understanding of videos in a dense and comprehensive manner, such as robotics, autonomous driving, and augmented-reality applications.
Metrics.
We evaluate our method using standard evaluation metrics in image instance segmentation, adapted for video instance segmentation [38]. These include Average Precision (AP) and Average Recall (AR) IoU-based metrics. Given that each instance in a video comprises a sequence of masks, unlike image instance segmentation, IoU computation is carried out not only in the spatial dimensions but also in the temporal dimension. This implies that the sum of intersections at every single frame is divided by the sum of unions at every single frame. These metrics are generally computed on a per-category basis and subsequently averaged across all categories. However, we work with the class-agnostic version of UVO.
4.2 Implementation Details
Training Data.
For our experiments, we use pre-trained checkpoints provided by the respective authors for both PIPS [11] and SAM. PIPS is trained exclusively on a synthetic dataset, FlyingThings++ [11], derived from the FlyingThings [24] optical flow dataset. This dataset includes multi-frame amodal trajectories with synthetic occlusions caused by moving objects. SAM, on the other hand, has been trained on the large-scale SA-1B dataset, the largest image segmentation dataset to date, with over 1 billion masks on 11M licensed and privacy-respecting images. It is noteworthy that neither of these datasets includes video segmentation data, and they do not overlap with any of our evaluation data. This effectively positions our model in a zero-shot video segmentation setting.
Model Variations.
Our experiments led to two optimal model hyperparameters, distinguished as SAM-PT (without reinitialization) and SAM-PT-reinit (with reinitialization). These configurations were derived from our ablation study in Sec. 4.3. However, we found that using iterative refinement negatively impacted both SAM-PT and SAM-PT-reinit on the MOSE dataset, and likewise hindered SAM-PT-reinit on the YouTube-VOS dataset. Consequently, iterative refinement was deactivated for these specific datasets. For DAVIS, we additionally report results for replacing SAM with HQ-SAM [15] and denote the model variants as HQ-SAM-PT and HQ-SAM-PT-reinit. The HQ-SAM variants use 3 iterative refinement iterations instead of 12 iterations.
VOS Evaluation.
When evaluating on VOS, we use the provided ground truth mask for the first frame to sample the query points required by our method. Then, we give only the sampled points as input to our method, not the mask. For all datasets, we use the full-resolution data and resize it to the longest side of 1024 to match SAM’s input resolution.
VIS Evaluation.
For evaluating our method on the VIS task, we leverage SAM’s automatic mask generation capacity to generate up to 100 mask proposals for the initial frame. We then propagate these proposed masks throughout the entire video sequence using our method. We evaluate TAM [40], a concurrent method we compare against, in the same manner. Our mask proposal generation process is currently simplistic and does not create any proposals for subsequent video frames. Consequently, it cannot identify objects that emerge in later frames, placing it at a disadvantage compared to VIS methods that are capable of doing so. Despite this limitation, our approach provides a consistent platform for comparing zero-shot methods in terms of how effectively they propagate diverse mask proposals from the first frame.
4.3 Ablation Study
We conducted detailed ablation experiments on the DAVIS $2017$ validation subset to validate various components and designs of SAM-PT. We employed SAM’s ViT-H as the backbone, for all tests. Each aspect was examined sequentially, integrating the optimal settings obtained from prior experiments. To ensure statistical soundness, multiple iterations of each experiment were carried out (between 4 and 12 runs per setup), with findings represented as mean and standard deviation across these runs.
While these results provide insight, there may be a risk of overfitting due to our limited validation dataset. While we endeavored to maintain a consistent evaluation protocol, future research should aim for a larger validation set, possibly derived from the YouTube VOS $2018$ train dataset, to mitigate this concern.
Figure 5: Query Point Sampling Ablation Study. We report mean $\mathcal{J\&F}$ scores along with the standard deviation on the validation subset of DAVIS $2017$ for different number of positive points per mask and different point selection methods. Using $8$ positive points per mask leads to a $40$-point performance boost compared to using a single point. Given $8$ positive points, K-Medoids and Shi-Tomasi perform comparably well as point selection methods.
Query Point Sampling
Fig. 5 illustrates that the number of positive points and the choice of point selection methods significantly influence performance. Using 8 points per mask showed a remarkable 40-point performance enhancement compared to a single point. This substantiates the argument that a single positive point is inadequate for prompting SAM as it often results in the segmentation of partial objects only. Among the point selection methods, K-Medoids and Shi-Tomasi produced comparable results, with a slight preference towards K-Medoids owing to its marginally higher mean score and resilience to the number of positive points per mask.
Table 2: Point Tracker and SAM Configuration Ablation Study Results. Using the best parameters from the ablation study in Fig. 5, we report the mean performance (with standard deviation) on the validation subset of DAVIS $2017$ to study the impact of (a) different point trackers, (b) the number of negative points per mask, (c) the use of iterative refinement and (d) patch similarity filtering. We find that the best configuration uses PIPS as the point tracker, $1$ negative point per mask, $12$ refinement iterations, and no patch similarity filtering. PT: point tracker. NP: negative points per mask. IRI: iterative refinement iterations. PS: point similarity filtering threshold.
SAM-PT Configuration DAVIS $\mathbf{2017}$ Validation [27]
[…]
(a) point tracker
RAFT [30] 00 00 $63.0\pm 0.6$ $60.7\pm 0.6$ $65.4\pm 0.5$
SuperGlue [28] 00 00 $21.7\pm 2.8$ $19.6\pm 2.1$ $23.8\pm 3.4$
SuperGlue [28] 00 $3$ $28.4\pm 3.1$ $24.7\pm 2.4$ $32.0\pm 3.8$
TapNet [9] 00 00 $60.9\pm 0.2$ $58.2\pm 0.3$ $63.5\pm 0.2$
PIPS [11] 00 00 $\mathbf{72.3}\pm 1.2$ $\mathbf{70.4}\pm 1.3$ $\mathbf{74.3}\pm 1.1$ $\mathbf{+9.3}$
(b) negative points per mask
PIPS 00 00 $72.3\pm 1.2$ $70.4\pm 1.3$ $74.3\pm 1.1$
PIPS $1$ 00 $\mathbf{74.1}\pm 0.7$ $\mathbf{72.1}\pm 0.6$ $\mathbf{76.1}\pm 0.7$ $\mathbf{+1.8}$
PIPS $8$ 00 $74.0\pm 0.8$ $71.9\pm 0.8$ $76.0\pm 0.9$
PIPS $16$ 00 $73.4\pm 0.6$ $71.4\pm 0.6$ $75.3\pm 0.6$
PIPS $72$ 00 $72.2\pm 0.4$ $70.3\pm 0.4$ $74.0\pm 0.4$
(c) iterative refinement iterations
PIPS $1$ 00 $74.1\pm 0.7$ $72.1\pm 0.6$ $76.1\pm 0.7$
PIPS $1$ $1$ $75.7\pm 0.7$ $73.4\pm 0.7$ $78.1\pm 0.6$
PIPS $1$ $3$ $76.0\pm 0.6$ $73.4\pm 0.7$ $78.6\pm 0.7$
PIPS $1$ $12$ $\mathbf{76.3}\pm 0.6$ $\mathbf{73.6}\pm 0.6$ $\mathbf{78.9}\pm 0.6$ $\mathbf{+2.2}$
(d) patch similarity filtering
PIPS $1$ $12$ $\mathbf{76.3}\pm 0.6$ $\mathbf{73.6}\pm 0.6$ $\mathbf{78.9}\pm 0.6$ none
PIPS $1$ $12$ $0.002$ $72.7\pm 2.0$ $70.2\pm 1.8$ $75.2\pm 2.1$
PIPS $1$ $12$ $0.01$ $70.7\pm 2.0$ $68.3\pm 1.8$ $73.2\pm 2.1$
Point Tracking.
Tab. 2a shows that PIPS [11] demonstrated superior performance over TapNet [9], SuperGlue [28], and RAFT [30]. TapNet’s limitations stem from its lack of effective time consistency and its training on 256x256 images, which hampered its performance with higher-resolution images. SuperGlue, while proficient in matching sparse features across rigid scenes, grapples with effectively matching points from the reference frame in dynamic scenes, particularly under object deformations. RAFT, being an optical flow model, faced difficulties handling occlusions. Although PIPS’s prior use in our experiments may have offered some hyperparameter advantages, its superior performance is primarily attributable to its more robust design that emphasizes trajectory modeling over eight subsequent frames. This approach fosters the generation of coherent point trajectories and enhances occlusion detection.
Negative Points.
Tab. 2b highlights that incorporating negative points had a favorable impact, particularly in reducing segmentation errors when points deviated from the target object. The addition of negative points empowered SAM to better handle the point trackers’ failure cases, leading to improved segmentation and a 1.8-point enhancement over the non-use of negative points. Note that throughout all experiments, we always used the mixed point sampling method for sampling negative points which amounts to using random sampling when there is only one negative point per mask.
Iterative Refinement.
The iterative refinement approach contributed to higher-quality masks and mitigated the impact of artifacts in SAM’s output. Tab. 2c displays that this yielded an improvement of 2.2 points over the non-refinement approach.
Patch Similarity.
Our initial findings in Tab. 2d suggest that using patch similarity to filter unreliable tracking points was overly restrictive in our context, leading to substantial deletion of points. Although it did not prove beneficial in our current setup, this aspect certainly warrants further exploration, particularly in scenarios involving point re-initialization.
Reinitialization.
Fig. 6 presents the performance of different reinitialization variants. In Tab. 5 and Tab. 6, we also show it brings 2.5 and 2.0 points improvements on MOSE and UVO benchmarks respectively. The re-initialization process enhanced robustness against points falling off objects. By reinitializing all points based on the current mask prediction, we account for errors in point tracker outputs by discarding incorrect points and starting fresh from the current mask prediction. However, this assumes that we trust the currently outputted mask, which may not always be the case and sometimes leads to failures.
Figure 6: Reinitialization Ablation Study. Mean $\mathcal{J\&F}$ scores (with std. dev.) on the validation subset of DAVIS $2017$ for different reinitialization variants and configurations. The best result is achieved by reinitialization variant $A$, 12 refinement iterations, and 72 negative points per mask. The reinitialization variants differ in how the timestep at which the points get reinitialized is chosen, see Appendix A for more details. Although using reinitialization improves the performance only marginally on the validation subset, this strategy demonstrates substantial improvement on the MOSE $2023$ (Tab. 5) and UVO (Tab. 6) datasets.
In summary, our best-performing SAM-PT model employs K-Medoids for point selection with 8 points per mask, PIPS for point tracking, a single negative point per mask, and employs 12 iterations for iterative refinement without patch similarity filtering. Meanwhile, using reinitialization achieved optimum performance with 12 refinement iterations and 72 negative points per mask.
4.4 Comparison with State-of-the-art Methods
All reported results were computed with official tools or official evaluation servers. Sec. 4.4.1 reports Video Object Segmentation results, including qualitative results on unseen web videos. Sec. 4.4.2 reports Video Instance Segmentation results.
4.4.1 Video Object Segmentation
Table 3: Quantitative results in semi-supervised VOS on the validation subset of DAVIS $2017$.
DAVIS $\mathbf{2017}$ Validation [27]
[…]
(a) trained on video segmentation data
AGSS [20] $67.4$ $64.9$ $69.9$
AGAME [13] $70.0$ $67.2$ $72.7$
AFB-URR [18] $74.6$ $73.0$ $76.1$
STM [25] $81.8$ $79.2$ $84.3$
SWEM [21] $84.3$ $81.2$ $87.4$
RDE [17] $86.1$ $82.1$ $90.0$
SwinB-DeAOT-L [41] $86.2$ $83.1$ $89.2$
XMem [4] $87.7$ $84.0$ $91.4$
(b) not trained on video segmentation data (zero-shot)
Painter [35] $34.6$ $28.5$ $40.8$
DINO [2] $71.4$ $67.9$ $74.9$
SegGPT [36] 75.675.6{75.6}75.6 72.572.5{72.5}72.5 78.678.6{78.6}78.6
PerSAM-F [45] $71.9$ $69.0$ $74.8$
SAM-PT (ours) $76.3\pm 0.6$ $73.6\pm 0.6$ ${78.9}\pm 0.6$
SAM-PT-reinit (ours) ${76.6}\pm 0.7$ ${74.4}\pm 0.8$ ${78.9}\pm 0.6$
HQ-SAM-PT [15] (ours) $\mathbf{77.2}\pm 0.5$ $74.7\pm 0.5$ $\mathbf{79.8}\pm 0.4$
HQ-SAM-PT-reinit [15] (ours) $77.0\pm 0.7$ $\mathbf{74.8}\pm 0.8$ $79.2\pm 0.6$
Performance Overview.
Our proposed method outperforms others that have not been trained on any video object segmentation data on the DAVIS $2017$ dataset, as reflected in Tab. 3. A mean $\mathcal{J\&F}$ score of $76.6$ points exceeds the PerSAM-F by $4.7$ points and the SegGPT generalist model by a single point. The experiments were repeated 8 times for statistical robustness, and we report the mean and standard deviation of our method’s performance.
We also outperform PerSAM-F on the YouTube-VOS $2018$ and MOSE $2023$ datasets, achieving mean scores of $67.0$ and $41.0$ as shown in Tabs. 4 and 5. However, with different mask training data, our performance falls short when compared to SegGPT on the two datasets.
(a) Good cases for SAM-PT
(b) Good cases for SAM-PT-reinit
(c) Failure cases for SAM-PT
Figure 7: Visualization of SAM-PT on the DAVIS 2017 Validation [27] dataset. Given the first-frame masks, we sample 8 positive points and either 1 (SAM-PT) or 72 (SAM-PT-reinit) negative points per object to initialize SAM-PT and SAM-PT-reinit. Circles denote positive points, crosses denote negative points, and red symbols (circle and cross) denote that the point was predicted to be occluded.
Figure 8: Successful segmentation using SAM-PT on short clips from “Avatar: The Last Airbender”. Although our method has never seen data from Avatar, an anime-influenced animated television series, it segments and tracks various objects in short clips.
Figure 9: Challenging scenarios for SAM-PT on short clips from “Avatar: The Last Airbender”. These cases illustrate instances where our model struggles when faced with point tracking failures that are the result of incorrectly predicting the point at a similar-looking segment or when faced with object occlusions and disappearing objects.
Qualitative Analysis.
Visualizations of successful video segmentation on DAVIS $2017$ for SAM-PT and SAM-PT-reinit can be seen in Fig. 6(a) and Fig. 6(b) respectively. Notably, Fig. 8 presents successful video segmentation on unseen web videos – clips from the “Avatar: The Last Airbender” anime-influenced animated television series, demonstrating the zero-shot capabilities of our method.
Limitations and Challenges.
Despite the competitive zero-shot performance, certain limitations persist, primarily due to the limitations of our point tracker in handling occlusion, small objects, motion blur, and re-identification. In such scenarios, the point tracker’s errors propagate into future video frames. Fig. 6(c) illustrates these problematic instances on DAVIS $2017$, while Fig. 9 presents additional cases on “Avatar: The Last Airbender” clips. Although using point re-initialization and negative points somewhat alleviates the failures of the point tracker, they still prevent the performance from being on par with methods trained on video data.
Table 4: Quantitative results in semi-supervised VOS on the validation subset of YouTube-VOS $2018$. Metrics are reported separately for “seen” and “unseen” classes, with $\mathcal{G}$ being the overall average score over the metrics. Note that SegGPT and SAM-PT adopt completely different training data. ♠: our reproduced result using the official code of [45].
YouTube-VOS $\mathbf{2018}$ Validation [38]
Method $\mathcal{G}$ $\mathcal{J}_{s}$ $\mathcal{F}_{s}$ $\mathcal{J}_{u}$ $\mathcal{F}_{u}$
(a) trained on video segmentation data
AGAME [13] $66.0$ $66.9$ $61.2$
AGSS [20] $71.3$ $71.3$ $65.5$ $75.2$ $73.1$
STM [25] $79.4$ $79.7$ $84.2$ $72.8$ $80.9$
AFB-URR [18] $79.6$ $78.8$ $83.1$ $74.1$ $82.6$
RDE [17] $83.3$ $81.9$ $86.3$ $78.0$ $86.9$
SWEM [21] $82.8$ $82.4$ $86.9$ $77.1$ $85.0$
XMem [4] $86.1$ $85.1$ $89.8$ $80.3$ $89.2$
SwinB-DeAOT-L [41] $86.2$ $85.6$ $90.6$ $80.0$ $88.4$
(b) not trained on video segmentation data (zero-shot)
Painter [35] $24.1$ $27.6$ $35.8$ $14.3$ $18.7$
SegGPT [36] $\mathbf{74.7}$ $\mathbf{75.1}$ $\mathbf{80.2}$ $\mathbf{67.4}$ $\mathbf{75.9}$
PerSAM-F ♠ [45] $54.4$ $53.9$ $56.4$ $50.7$ $56.6$
SAM-PT (ours) $67.0\pm 0.3$ $68.6\pm 0.2$ $71.2\pm 0.1$ $61.0\pm 0.5$ $67.4\pm 0.4$
SAM-PT-reinit (ours) ${67.5}\pm 0.2$ ${69.0}\pm 0.4$ ${69.9}\pm 0.3$ ${63.2}\pm 0.4$ ${67.8}\pm 0.5$
Table 5: Quantitative results in semi-supervised VOS on the validation subset of MOSE $2023$ [8]. ♠: our reproduced result using the official code of [45].
MOSE $\mathbf{2023}$ Validation [8]
[…]
(a) trained on video segmentation data
RDE [17] $48.8$ $44.6$ $52.9$
SWEM [21] $50.9$ $46.8$ $54.9$
XMem [4] $57.6$ $53.3$ $62.0$
DeAOT [41] $59.4$ $55.1$ $63.8$
(b) not trained on video segmentation data (zero-shot)
Painter [35] $14.5$ $10.4$ $18.5$
SegGPT [36] $\mathbf{45.1}$ $\mathbf{42.2}$ $\mathbf{48.0}$
PerSAM-F ♠ [45] $23.3$ $19.8$ $26.8$
SAM-PT (ours) $38.5\pm 0.2$ $34.9\pm 0.3$ $42.1\pm 0.2$
SAM-PT-reinit (ours) ${41.0}\pm 0.5$ ${38.5}\pm 0.5$ ${43.5}\pm 0.5$
4.4.2 Video Instance Segmentation
Results and Analysis.
Given the same mask proposals, SAM-PT outperforms TAM [40] significantly even though SAM-PT was not trained on any video segmentation data. TAM is a concurrent approach combining SAM and XMem [4], where XMem was pre-trained on BL30K [5] and trained on DAVIS and YouTube-VOS, but not on UVO. On the other hand, SAM-PT combines SAM with the PIPS point tracking method, both of which have not been trained on video segmentation tasks.
5 Conclusion
We present SAM-PT, an innovative solution that extends SAM’s segmentation ability from static images to dynamic videos. Integrated with long-term point trackers, our approach demonstrates strong performance across several benchmarks including DAVIS, YouTube-VOS, MOSE, and UVO. While our method has limitations such as difficulty handling occlusions, small objects, and motion blur, and inconsistencies in mask predictions, it contributes a simple and effective new point-based perspective to video object segmentation research. By illustrating a promising way to extend foundational models like SAM into the video domain, our research provides a potential pathway for advancements in diverse applications from autonomous driving to video labeling. Furthermore, the future incorporation of more advanced point trackers can enhance the performance of SAM-PT.
[…]
Method Propagation $\mathbf{AR100}$ $\mathbf{ARs}$ $\mathbf{ARm}$ $\mathbf{ARl}$ $\mathbf{AP}$
(a) trained on video segmentation data, including UVO’s training subset
Mask2Former VIS [44] N/A $35.4$- -- -- $27.3$
ROVIS [44] N/A $41.2$- -- -- $32.7$
(b) trained on video segmentation data
TAM [40] Feature Matching $24.1$ $21.1$ $32.9$ $31.1$ $1.7$
(c) not trained on video segmentation data (zero-shot)
SAM-PT (ours) Points Prompting $28.8\pm 0.1$ $23.3\pm 0.1$ $40.8\pm 0.2$ $48.3\pm 0.6$ $\mathbf{6.7}\pm 0.2$
SAM-PT-reinit (ours) Points Prompting $\mathbf{30.8}$ $\mathbf{25.1}$ $\mathbf{44.1}$ $\mathbf{49.2}$ $6.5$
Appendix A Point Tracking Reinitialization
In our SAM-PT-reinit method, we introduce a reinitialization strategy. Here, the point tracker begins anew after every $h$ frames, where $h$ represents a pre-set tracking horizon (e.g., 8 frames), or is dynamically determined based on SAM’s mask predictions for each timestep within the horizon (e.g., using most-similar-mask-area heuristics). Upon reaching this horizon, the query points given to the tracker are reinitialized according to the mask prediction SAM outputted at the horizon frame. While this method may increase the computational load (especially if some of SAM’s computed masks are disregarded), it demonstrates substantial performance improvement in demanding video sequences, such as those in the MOSE dataset.
We explored four reinitialization methods, each varying in how they compute the value of $h$:
  1. (A)
    Reinit-on-Horizon-and-Sync-Masks: This straightforward variant reinitializes points after a fixed number of frames (e.g., every 8 frames). However, it may stumble if the mask is absent at the reinitialization timestep. Despite this potential pitfall, it operates at the same speed as methods that do not employ reinitialization.
  2. (B)
    Reinit-at-Median-of-Area-Diff: In this variant, the tracker outputs trajectory points for each frame within the horizon, and SAM predicts masks based on these trajectories. Reinitialization happens at the frame within the horizon that has the mean mask area among the non-empty masks predicted by SAM. Notably, this approach may be significantly slower than methods without reinitialization, as it may reject several SAM masks (e.g., out of 8 computed masks, reinitialization might occur on the second one, necessitating recomputation of the remaining 6 masks in the next step).
  3. (C)
    Reinit-on-Similar-Mask-Area: This method triggers reinitialization when the mask area is similar to the initial mask area, causing it to be several times slower than methods without reinitialization.
  4. (D)
    Reinit-on-Similar-Mask-Area-and-Sync-Masks: This variant reinitializes when the mask area for all masks in the batch is similar to the initial mask areas, synchronizing the masks to be tracked from the same timestep. This synchronization allows for the use of negative points from other masks when querying SAM, but it also runs several times slower than methods without reinitialization.
From our investigations, we found the (A) Reinit-on-Horizon-and-Sync-Masks strategy to be the most effective, as indicated by its superior performance on the DAVIS $2017$ validation subset. The choice of reinitialization method may depend on the specific validation subset and the degree of hyperparameter tuning involved. Note that we have always used reinitialization along with negative points.
A.1 Computational Cost and Speed Optimization
The introduction of reinitialization in SAM-PT-reinit comes with a trade-off: it slows down the inference speed by a factor of 2 to 8, depending on the reinitialization method and parameters used. The major bottleneck is the invocation of SAM’s backbone for each video frame. We propose caching the backbone outputs for unprocessed video frames as a possible solution to mitigate this slowdown. This strategy requires storing embeddings for all video frames in the working memory but offers the potential for significant speedup, particularly useful for applications requiring faster inference.
Appendix B More DAVIS Subsets
We report results on DAVIS $2016$ Validation and DAVIS $2017$ Test-dev in Tab. 7.
Table 7: Quantitative results in semi-supervised video object segmentation on additional subsets of DAVIS. ♠: our reproduced result using the official code of [45].
DAVIS $\mathbf{2016}$ Validation [27] DAVIS $\mathbf{2017}$ Test-dev [27]
[…]
PerSAM-F ♠ [45] $74.8$ $74.5$ $75.0$ $47.6$ $45.5$ $49.7$
SegGPT [36] $82.3$ $81.8$ $82.8$- - -
SAM-PT $\mathbf{83.1}\pm 1.5$ $\mathbf{83.0}\pm 0.8$ $\mathbf{83.0}\pm 1.1$ $\mathbf{62.7}\pm 0.5$ $\mathbf{59.4}\pm 0.6$ $\mathbf{66.1}\pm 0.4$
SAM-PT-reinit $80.2\pm 0.6$ $80.3\pm 0.6$ $80.0\pm 0.6$ $61.5\pm 1.1$ $59.3\pm 1.0$ $63.8\pm 1.2$
Appendix C Per-sequence DAVIS 2017 Validation Results
See figure Fig. 10 for per-sequence DAVIS 2017 Validation results. For exact numbers and tables, check our GitHub experiment summaries or the Wandb project.
(a) number of positive points per mask
(b) point selection methods
(c) point tracking module
(d) number of negative points per mask
(e) number of iterative refinement iterations
(f) patch similarity filtering threshold
Figure 10: Per-sequence Ablation Results. The charts plot the median J-Mean result of different ablation result experiments along with $95\%$ bootstrapped confidence intervals, based on 1000 bootstrap samples.

### v2

@@
 Abstract
-The Segment Anything Model (SAM) has established itself as a powerful zero-shot image segmentation model, employing interactive prompts such as points to generate masks.
-This paper presents SAM-PT, a method extending SAM’s capability to tracking and segmenting anything in dynamic videos.
-SAM-PT leverages robust and sparse point selection and propagation techniques for mask generation, demonstrating that a SAM-based segmentation tracker can yield strong zero-shot performance across popular video object segmentation benchmarks, including DAVIS, YouTube-VOS, and MOSE.
-Compared to traditional object-centric mask propagation strategies, we uniquely use point propagation to exploit local structure information that is agnostic to object semantics.
+The Segment Anything Model (SAM) has established itself as a powerful zero-shot image segmentation model, enabled by efficient point-centric annotation and prompt-based models.
+While click and brush interactions are both well explored in interactive image segmentation, the existing methods on videos focus on mask annotation and propagation.
+This paper presents SAM-PT, a novel method for point-centric interactive video segmentation, empowered by SAM and long-term point tracking.
+SAM-PT leverages robust and sparse point selection and propagation techniques for mask generation.
+Compared to traditional object-centric mask propagation strategies, we uniquely use point propagation to exploit local structure information agnostic to object semantics.
 We highlight the merits of point-based tracking through direct evaluation on the zero-shot open-world Unidentified Video Objects (UVO) benchmark.
-To further enhance our approach, we utilize K-Medoids clustering for point initialization and track both positive and negative points to clearly distinguish the target object.
-We also employ multiple mask decoding passes for mask refinement and devise a point re-initialization strategy to improve tracking accuracy.
-Our code integrates different point trackers and video segmentation benchmarks and will be released at https://github.com/SysCV/sam-pt.
+Our experiments on popular video object segmentation and multi-object segmentation tracking benchmarks, including DAVIS, YouTube-VOS, and BDD100K, suggest that a point-based segmentation tracker yields better zero-shot performance and efficient interactions.
+We release our code that integrates different point trackers and video segmentation benchmarks at https://github.com/SysCV/sam-pt.
 {strip}
 Figure 1: Segment Anything Meets Point Tracking (SAM-PT).
-SAM-PT is the first method to utilize sparse point propagation for Video Object Segmentation (VOS).
-The essence of SAM-PT is to extend SAM [16] with long-term point trackers to effectively operate on videos in a zero-shot manner.
-SAM-PT takes a video as input together with annotations of the target object in the first frame.
-These annotations are called “query points” and denote either the target object (positive points) or designate non-target segments (negative points).
-The points are tracked throughout the video using point trackers that propagate the query points to all video frames, producing predicted trajectories and occlusion scores.
+SAM-PT is a point-centric method that utilizes sparse point propagation for interactive video segmentation, enabling easier interaction and faster annotation.
+We extend SAM [21] with long-term point trackers to effectively operate on videos in a zero-shot manner.
+SAM-PT takes user clicks as “query points” which either denote the target object (positive points) or designate non-target segments (negative points).
+The points are tracked throughout the video using point trackers that propagate the query points to all video frames, producing trajectory predictions and occlusion scores.
 SAM is subsequently prompted with the non-occluded points in the trajectories as to output a segmentation mask for each video frame independently.
+The propagated points can be further edited for accurate segmentation and tracking.
 1 Introduction
-Video segmentation benefits a myriad of applications, including autonomous driving, robotics, and video editing.
-Despite significant progress made in the past few years with deep neural networks [41, 4, 37, 3], the current methodologies falter when faced with unseen data, particularly in zero-shot settings.
-These models struggle to maintain consistent performance across diverse scenarios without specific video segmentation data for fine-tuning.
-The prevailing methods [4, 3] in semi-supervised Video Object Segmentation (VOS) and Video Instance Segmentation (VIS) exhibit performance gaps when dealing with unseen data, particularly in a zero-shot setting, i.e., when these models are transferred to video domains they have not been trained on and encompass object categories that fall outside of the training distribution.
-A potential route towards overcoming these challenges lies in adapting successful models in the image segmentation domain for video segmentation tasks.
-One such promising model is the Segment Anything Model (SAM) [16].
-SAM is a powerful foundation model for image segmentation, trained on the large-scale SA-1B dataset, which contains an astounding 11 million images and over 1 billion masks.
-This extensive training set enables SAM’s impressive zero-shot generalization capabilities.
-The model is highly adaptable, able to produce high-quality masks from single foreground points, and has demonstrated robust performance across a range of downstream tasks under zero-shot transfer protocols.
-While SAM demonstrates powerful zero-shot capabilities for image segmentation, it is not innately suited for video segmentation tasks.
-Recent efforts have been made to adapt SAM for video segmentation.
-For instance, TAM [40] integrates SAM with the state-of-the-art memory-based mask tracker XMem [4].
-Likewise, SAM-Track [6] combines SAM with DeAOT [41].
-While these methods mostly recover the performance on in-distribution data, they fall short in preserving the original performance of SAM in more challenging, zero-shot settings.
-Other methods that do not leverage SAM, such as SegGPT [36], can successfully solve a number of segmentation problems using visual prompting, but still require mask annotation for the first video frame.
-This problem represents a significant barrier in zero-shot video segmentation, particularly as we seek to develop methods that can easily generalize to unseen scenarios and consistently deliver high-quality segmentation across diverse video domains.
-We introduce SAM-PT (Segment Anything Meets Point Tracking), depicted in Segment Anything Meets Point Tracking.
+Object segmentation and tracking in videos are central pillars for a myriad of applications, including autonomous driving, robotics, and video editing.
+Despite significant progress made in the past few years with deep neural networks [48, 7, 43, 5], we still need to rely on expensive labels for model supervision to achieve high accuracies.
+Therefore, many efforts have been made on interactive video segmentation to accelerate the data labeling process and benefit artistic video editing.
+Those methods are usually evaluated on a simplified semi-supervised segmentation setup as a proxy for full interactive video segmentation.
+The prevailing methods [7, 5] in semi-supervised Video Object Segmentation (VOS) and Video Instance Segmentation (VIS) exhibit performance gaps when dealing with unseen data, particularly in a zero-shot setting, i.e., when these models are transferred to video domains they have not been trained or that encompass object categories falling outside of the training distribution.
+Further, generalizable models usually require large amounts of training data.
+The existing interactive video segmentation methods assume the mask of an object is given on the first frame of the testing video.
+While getting accurate mask is laborious, recent works [21, 6] on training foundation image segmentation models show that point-based annotation in combination with mask editing tools is a scalable approach to label exceedingly large amounts of data.
+Despite its success on images in terms of labeling efficiency and accuracy, point-centric interactive segmentation has received scant attention in the video domains.
+In this paper, we aim to achieve both domain generalizability and labeling efficiency for interactive video segmentation.
+Our insights are two-fold.
+First, foundation models in image segmentation are available, such as Segment Anything Model (SAM) [21].
+SAM, trained on 11 million images and 1 billion object masks, has impressive zero-shot generalization capabilities.
+The model also supports point prompts as additional inputs for interactive image segmentation and produces high-quality masks.
+Second, we witnessed significant recent progress in point tracking [14, 16, 53, 15, 39, 19].
+Those tracking methods, once trained, can propagate points across video frames on diverse domains.
+Therefore, we introduce SAM-PT (Segment Anything Meets Point Tracking), depicted in Segment Anything Meets Point Tracking.
 This is the first method to utilize sparse point tracking combined with SAM for video segmentation, offering a new perspective on solving the problem.
-Instead of employing object-centric dense feature matching or mask propagation, we propose a point-driven approach that capitalizes on tracking points using rich local structure information embedded in videos.
-As a result, it only requires sparse points annotation to denote target object in the first frame and provides better generalization to unseen objects, a strength demonstrated on the open-world UVO [34] benchmark.
+Instead of employing object-centric dense feature matching or mask propagation, we propose a point-centric approach that capitalizes on tracking points using rich local structure information embedded in videos.
+It only requires sparse points annotation to denote the target object in the first frame and provides better generalization to unseen objects.
 This approach also helps preserve the inherent flexibility of SAM while extending its capabilities effectively to video segmentation.
-SAM-PT prompts SAM with sparse point trajectories predicted using state-of-the-art point trackers, such as PIPS [11], harnessing their versatility for video segmentation.
-We identify that initializing points to track using K-Medoids cluster centers from a mask label was the strategy most compatible with prompting SAM.
+Similar to the data annotation process in SAM, our point-centric approach can be potentially integrated with the existing mask-based approaches in real-world applications.
+SAM-PT prompts SAM with sparse point trajectories predicted using state-of-the-art point trackers, such as CoTracker [19], harnessing their versatility for video segmentation.
+We identified that initializing points to track using K-Medoids cluster centers from a mask label was the strategy most compatible with prompting SAM.
 Tracking both positive and negative points enables the clear delineation of target objects from their background.
 To further refine the output masks, we propose multiple mask decoding passes that integrate both types of points.
-In addition, we devised a point re-initialization strategy that increases tracking accuracy over time.
+In addition, we devised a point reinitialization strategy that increases tracking accuracy over time.
 This approach involves discarding points that have become unreliable or occluded, and adding points from object parts or segments that become visible in later frames, such as when the object rotates.
-Notably, our experimental results highlight that SAM-PT competes with existing zero-shot methods [36] or outperforms them [45, 2, 12, 35] on several video segmentation benchmarks.
-This comes without the need for any video segmentation data during training, underscoring the robustness and adaptability of our approach.
-SAM-PT holds the potential to enhance progress in video segmentation tasks, particularly in zero-shot scenarios.
+We evaluate SAM-PT on multiple setups including semi-supervised, open-world, and fully interactive video segmentation.
+Our method achieves stronger performance than existing zero-shot methods by up to $5.0\%$ on DAVIS, $2.0\%$ on YouTube-VOS, and $7.3\%$ on BDD100K, while also surpassing a fully-supervised VIS method [46] on UVO by $6.7$ points.
+We also set up a new benchmark for interactive point-based video segmentation to simulate the process of manually labeling the whole video.
+In this setup, SAM-PT significantly reduces annotation effort, approaching the performance of fully supervised approaches and underscoring its practicality.
+This comes without the need for any video segmentation data during training, underscoring the robustness and adaptability of our approach, and indicating its potential to enhance progress in video segmentation tasks, particularly in zero-shot scenarios.
 2 Related Work
@@
 Classical feature extraction and tracking methods such as Lucas-Kanade [26], Tomasi-Kanade [36], Shi-Tomasi [34], SIFT [25], and SURF [1], as well as newer methods such as LIFT [49], SuperPoint [12], and SuperGlue [33], have all demonstrated proficiency in identifying or tracking sparse features and establishing long-range correspondences.
-However, their effectiveness is confined to a specific set of distinct interest points and they often struggle when applied to non-rigid, dynamic scenes.
-Flow-based methods, such as RAFT [30], excel in tracking dense points between successive frames.
-However, they stumble with deriving accurate long-range point trajectories.
-When chaining flow predictions over time, errors tend to accumulate and lead to drift, while occlusions result in tracking failures.
-Significant strides have recently been made in long-term point tracking across video frames, as evinced by methods such as TapNet [9] and PIPS [11], as well as the concurrent and state-of-the-art OmniMotion [32] and TAPIR [10] techniques.
-These approaches optimize long-range point trajectories across an entire video, navigating mostly well through periods of occlusion.
-Our work stands apart as the first to integrate these successful long-term point tracking methods, utilizing them to guide a promptable foundation model for image segmentation toward performing video segmentation tasks.
+Nonetheless, these techniques often falter in dynamic, non-rigid environments.
+While flow-based approaches such as RAFT [35] offer improvements, they too struggle with maintaining long-term point accuracy due to error accumulation and occlusions.
+Addressing these shortcomings, recent innovations such as PIPS [16], PIPS++ [53], OmniMotion [39], TAPIR [15], and the state-of-the-art CoTracker [19], optimize for robust long-term trajectories and effectively manage occlusions.
+Our work is unique in applying these methods to guide image segmentation models for video segmentation tasks.
 Segment and Track Anything Models.
-SAM [16] is an innovative image segmentation model for promptable image segmentation, trained on over 1 billion segmentation masks.
-It showcases remarkable zero-shot generalization abilities and can produce high-quality masks from a single foreground point.
-To further improve the quality of the masks, especially when segmenting objects with intricate structures, HQ-SAM [15] extends SAM with a learnable high-quality output token which proves efficient in diverse segmentation domains.
-However, SAM and HQ-SAM cannot be directly used to solve video segmentation tasks.
-A few concurrent works extend SAM, for example, TAM [40] and SAM-Track [6] combine SAM with state-of-the-art mask trackers (such as XMem [4] and DeAOT [42]) to perform interactive video object segmentation.
-These methods employ SAM for mask initialization or correction and XMem/DeAOT for mask tracking and prediction.
-Using the pre-trained mask trackers recovers the in-distribution performance, but hinders the performance in zero-shot settings.
-PerSAM [45] also demonstrates the ability to track multiple reference objects in a video.
-Instead of building an interactive tracking pipeline or SAM fine-tuning, we focus on learning robust associations for diverse objects in zero-shot scenarios.
+SAM [21] is a foundation model for image segmentation that showcases impressive zero-shot capabilities.
+Its extension, HQ-SAM [20], improves mask quality for complex objects but is not designed for video tasks.
+TAM [46] and SAM-Track [11] attempt to extend SAM to video segmentation by integrating the state-of-the-art fully-supervised XMem [7] and DeAOT [47] mask trackers, respectively, yet they lack in zero-shot scenarios.
 Zero-Shot VOS / VIS.
-Among the non-SAM-based methods, Painter [35] and its SegGPT [36] extension are another sort of generalist models for solving a variety of image and segmentation tasks.
-These methods likewise use visual prompting techniques but are inherently different frameworks from SAM.
-Despite its wide applicability, Painter shows lacking performance in video segmentation tasks.
-Conversely, SegGPT successfully uses in-context prompting to achieve one-shot video object segmentation performance comparable to ours, also without training on any video data.
-The training domains, however, notably differ between SegGPT and our method.
-STC [12] and DINO [2] also do not use any video segmentation data during training.
-In the semi-supervised video object segmentation, they take a reference mask as input and perform frame-by-frame feature matching, which propagates the reference mask across the entirety of the video.
-Our SAM-PT, on the other hand, diverges substantially from these methodologies by adopting point tracking, eschewing the process of frame-by-frame feature matching.
-Additionally, our method requires only sparse points to represent the target object, rather than a full reference mask, and yields superior performance on conventional semi-supervised video object segmentation benchmarks.
+Generalist models such as Painter [41] apply visual prompting to various tasks but demonstrate limited performance in video segmentation.
+On the other hand, SegGPT [42] also uses visual prompting and competes closely with our method on some datasets.
+Other approaches, such as STC [18] and DINO [4], perform VOS through feature matching.
+Our approach distinguishes itself by taking the point-centric approach to enhance performance on VOS benchmarks in a zero-shot setting.
+Interactive VOS.
+Interactive VOS has shifted from labor-intensive manual annotations to more user-friendly interaction methods, such as scribbles, clicks, and drawings, enabling rapid and intuitive video editing [37, 28, 27, 3, 17].
+Among these, MiVOS [8] stands out for its modular design that decouples mask generation from propagation, effectively incorporating user interactions from diverse interaction modalities.
+Unlike MiVOS and other fully-supervised methods, SAM-PT is the first to use point propagation instead of mask propagation and thus operates effectively in zero-shot settings.
+Our interactive point-based video segmentation study emphasizes the simplicity and efficacy of point interactions and differs from common scribble-based benchmarking [2] or in-distribution user studies [8].
 3 Method
-We propose SAM-PT to adapt SAM, a foundation model for image segmentation, for addressing video segmentation tasks in a zero-shot setting.
-SAM-PT combines the strengths of existing prominent point trackers, such as PIPS [11] and TapNet [9], with the powerful image segmentation of SAM to enable tracking of anything in videos.
-First, Sec.
-3.1 briefly describes SAM.
+We propose SAM-PT for addressing video segmentation tasks in a zero-shot setting.
+SAM-PT combines the strengths of the Segment Anything Model (SAM), a foundation model for image segmentation, and prominent point trackers, such as PIPS [16] and CoTracker [19], to enable interactive tracking of anything in videos.
+Sec.
+3.1 briefly describes the background knowledge about SAM.
 Sec.
@@
 Finally, Sec.
-3.3 analyzes and highlights the method’s novelty as the first point-driven video segmentation method compared to existing works.
+3.3 analyzes and highlights the method’s novelty as the first point-centric interactive video segmentation method compared to existing works.
 3.1 Preliminaries: SAM
+Whereas in computer vision “zero-shot (learning)” usually refers to the study of generalization to unseen object categories in image classification [22], we follow prior work [21, 31] and rather employ the term in a broader sense and explore generalization to unseen datasets.
 The Segment Anything Model (SAM) [21] is a novel vision foundation model designed for promptable image segmentation.
 SAM is trained on the large-scale SA-1B dataset, which contains 11 million images and over 1 billion masks.
-SA-1B has 400 times more masks than any existing segmentation dataset.
-This extensive training set facilitates SAM’s impressive zero-shot generalization capabilities to new data.
-SAM has showcased its ability to produce high-quality masks from a single foreground point, and has demonstrated robust generalization capacity on a variety of downstream tasks under a zero-shot transfer protocol using prompt engineering.
+SA-1B has 400 times more masks than any prior segmentation dataset.
+This extensive training set facilitates SAM’s impressive zero-shot generalization capabilities.
+SAM has showcased its ability to produce high-quality masks from a single foreground point and has demonstrated robust generalization capacity on a variety of downstream tasks under a zero-shot transfer protocol using prompt engineering.
 These tasks include, but are not limited to, edge detection, object proposal generation, and instance segmentation.
-SAM comprises of three main components: an image encoder, a flexible prompt encoder, and a fast mask decoder.
+SAM comprises three main components: an image encoder, a flexible prompt encoder, and a fast mask decoder.
 The image encoder is a Vision Transformer (ViT) backbone and processes high-resolution $1024\times 1024$ images to generate an image embedding of $64\times 64$ spatial size.
@@
 Figure 2: Segment Anything Meets Point Tracking (SAM-PT) overview.
-The essence of SAM-PT is to extend image segmentation foundation model to effectively operate on videos.
+The essence of SAM-PT is to extend image segmentation foundation models to effectively operate on videos.
 SAM-PT has four steps: $\mathbf{1}$) Query Points Selection.
@@
 Optionally, the predicted masks are used to reinitialize the query points and restart the process when reaching a prediction horizon $h$.
-Re-initialization helps by getting rid of unreliable and occluded points, and adds points from object parts or segments that become visible in later frames, such as when the object rotates.
+Reinitialization helps by getting rid of unreliable points and adding points to object segments that become visible in later frames.
 1) Query Points Selection.
@@
   • Mixed Sampling: A hybrid method combining the above techniques since it might benefit from the unique strengths of each.
-While each method contributes distinct characteristics that influence the model’s performance, our ablation study reveals that K-Medoids sampling yields the best results with a good full coverage of various segments of the complete object.
+While each method contributes distinct characteristics that influence the model’s performance, our ablation study reveals that K-Medoids sampling yields the best results with good coverage of various segments of the complete object.
 Shi-Tomasi sampling follows closely, indicating their respective strengths in this context.
@@
 Initiated with the query points, we employ robust point trackers to propagate the points across all frames in the video, resulting in point trajectories and occlusion scores.
-We adopt the state-of-the-art point tracker PIPS [11] to propagate the points as PIPS shows moderate robustness toward long-term tracking challenges such as object occlusion and re-appearance.
-This is also shown more effective than methods such as chained optical flow propagation or first-frame correspondences in our experiment section.
+We adopt point trackers such as PIPS [16] and the state-of-the-art CoTracker [19] to propagate the points as they show moderate robustness toward long-term tracking challenges such as object occlusion and re-appearance.
+Long-term point trackers are also shown more effective than methods such as chained optical flow propagation or first-frame correspondences in our experiments.
 3) Segmentation.
@@
 4) Point Tracking Reinitialization.
-We optionally execute a reinitialization of the query points using the predicted masks once a prediction horizon of $h=8$ frames is reached, and denote the variant as SAM-PT-reinit.
-Upon reaching this horizon, we have $h$ predicted masks and will take the last predicted mask to sample new points.
+We optionally execute a reinitialization of the query points using the predicted masks once a prediction horizon of $h=8$ frames is reached.
+Upon reaching this horizon, we have $h$ predicted masks and will take the last one to sample new points.
 At this stage, all previous points are discarded and substituted with the newly sampled points.
@@
 The steps are iteratively executed until the entire video is processed.
-The reinitialization process serves to enhance tracking accuracy over time by discarding points that have become unreliable or occluded, while incorporating points from object segments that become visible later in the video.
-Other reinitialization variants are discussed in Appendix A and included in the ablation study in Sec.
-4.3.
+The reinitialization process serves to enhance tracking accuracy over time by discarding unreliable or occluded points while incorporating points from object segments that become visible later in the video.
+Other reinitialization variants are discussed in our Supplementary Material.
 3.3 SAM-PT vs. Object-centric Mask Propagation
@@
 1.
-To propagate the first-frame GT label to the remaining video frames, traditional techniques commonly use feature matching with masks cached to a mask memory [40, 6, 4, 41], frame-by-frame feature matching [12, 2], feature matching with the first-frame mask [45], optical flow [39], and, recently, in-context visual prompting [35, 36].
+To propagate the first-frame GT label to the remaining video frames, traditional techniques commonly use feature matching with masks cached to a mask memory [46, 11, 7, 48], frame-by-frame feature matching [18, 4], optical flow [45], and, recently, in-context visual prompting [41, 42].
 In contrast, SAM-PT introduces a unique approach to video object segmentation, employing the robust combination of point tracking with SAM, which is inherently designed to operate on sparse point prompts.
@@
 Further, the potential of our model extends beyond video object segmentation to other tasks, such as Video Instance Segmentation (VIS), thanks to the inherent flexibility of our point propagation strategy.
-In summary, SAM-PT is the first method that introduces sparse point propagation combined with prompting a image segmentation foundation model to perform zero-shot video object segmentation.
-It provides a fresh perspective and adds a new dimension to the study of video object segmentation.
-Table 1: Comparison of semi-supervised Video Object Segmentation (VOS) methods with respect to mask annotation requirements and propagation techniques.
-Our method, SAM-PT, is the first method for VOS that uses sparse point propagation.
-With such compact mask representation, we achieve the highest $\mathcal{J\&F}$ scores on the DAVIS $2016$ and $2017$ validation subsets among methods that do not utilize any video segmentation data during training.
-The methods are compared based on their use of video mask data during training, whether they are evaluated zero-shot on DAVIS, what first-frame labels they require, and what label propagation technique they employ.
-Method Video Mask Zero- Shot Frame Init. Propagation DAVIS $\mathbf{2016}$ DAVIS $\mathbf{2017}$
+Table 1: Comparative analysis of semi-supervised Video Object Segmentation methods.
+Our approach, SAM-PT, introduces sparse point propagation, a compact mask representation that uses local structure information agnostic to object semantics.
+It outperforms other non-video-data-dependent methods, achieving top $\mathcal{J\&F}$ scores on DAVIS $2016$ and $2017$, and the highest $\mathcal{G}$ score on YouTube-VOS $2018$.
+The comparison considers the reliance on video mask data during training, zero-shot learning setting, initial frame label requirements, and label propagation techniques used.
+Method Video Mask Zero- Shot Frame Init. Propagation DAVIS $\mathbf{2016}$ DAVIS $\mathbf{2017}$ YTVOS $\mathbf{2018}$
 SiamMask [38] Box Feature Correlation $69.8$ $56.4$
@@
 SAM-Track [11] Points Feature Matching $92.0$
-XMem [4] Mask Feature Matching $92.0$ 87.7
-DeAOT [41] Mask Feature Matching $\mathbf{92.9}$ $86.2$
-Painter [35] Mask Mask Prompting - $34.6$
+DEVA [10] Mask Feature Matching - $87.6$
+XMem [7] Mask Feature Matching $92.0$ 87.7 $86.1$
+DeAOT [48] Mask Feature Matching $\mathbf{92.9}$ $86.2$ $\mathbf{86.2}$
+Painter [41] Mask Mask Prompting - $34.6$ $24.1$
 STC [18] Mask Feature Matching - $67.6$
 DINO [4] Mask Feature Matching - $71.4$
-PerSAM-F [45] Mask Feature Matching - $71.9$
-SegGPT [36] Mask Mask Prompting $82.3$ $75.6$
-SAM-PT (ours) Points Points Prompting $\mathbf{83.1}$ $\mathbf{76.6}$
+SegGPT [42] Mask Mask Prompting $82.3$ $75.6$ $74.7$
+SAM-PT (ours) Points Points Prompting $\mathbf{84.3}$ $\mathbf{79.4}$ $\mathbf{76.2}$
 4 Experiments
 4.1 Datasets
-In the following subsections, we present an overview of the datasets used in our study.
-Section 4.1.1 provides a brief introduction to the Video Object Segmentation task and outlines the specific datasets we utilize for this task.
-Similarly, Section 4.1.2 discusses the Video Instance Segmentation task and the dataset associated with it.
-4.1.1 Video Object Segmentation
-Video Object Segmentation (VOS) refers to the process of segmenting a specific object across an entire video sequence.
-Semi-supervised VOS (also known as one-shot VOS or semi-automatic VOS) is the primary setting for VOS on which we evaluate our method.
-In this setting, the ground truth object mask of the first frame is provided, and the task is to predict the masks for subsequent frames.
-Alternatively, the first frame label can be a bounding box instead of a segmentation mask, or a set of points as is the case for our method.
 We evaluate our method on four VOS datasets: DAVIS $2016$, DAVIS $2017$ [30], YouTube-VOS $2018$ [44], and MOSE $2023$ [13].
-DAVIS $\mathbf{2016}$ [27].
-DAVIS $2016$ is a single-object VOS benchmark, consisting of $20$ highly diverse video sequences, each of which possesses well-annotated segmentation masks.
-DAVIS $\mathbf{2017}$ [27].
-A multi-object extension of its $2016$ version, DAVIS $2017$ includes $60$ videos in the training set and $30$ videos in the validation set, comprising a total of $197$ different objects.
-The video scenarios within this dataset are small but diverse.
-YouTube-VOS $\mathbf{2018}$ [38].
-YouTube-VOS $2018$ is a large-scale dataset collected from YouTube, comprising $3471$ training videos encompassing $65$ categories and $474$ validation videos with an additional $26$ unseen categories.
-The diversity in categories and the inclusion of seen and unseen classes allow for a comprehensive evaluation of a given model’s generalization capability.
-MOSE $\mathbf{2023}$ [8].
-MOSE $2023$ is a recently introduced dataset designed for multiple object segmentation and tracking in complex scenes.
-This dataset is replete with challenges such as the transient visibility of objects, the presence of minute or less noticeable entities, extensive occlusions, and scenes with a high object density.
-By design, each video in this dataset must contain multiple objects so that occlusions must be present, and objects must show sufficient motion, as opposed to being stationary or showing little movement.
-Metrics.
-We report the standard evaluation metrics for video object segmentation [8, 27, 38], including region similarity $\mathcal{J}$, contour accuracy $\mathcal{F}$, and their average, $\mathcal{J\&F}$.
-4.1.2 Video Instance Segmentation
-Video Instance Segmentation (VIS) is a task that combines object detection, instance segmentation, and object tracking across video frames, which aims to identify and segment each object instance over the whole video sequence.
-This is a much less explored task compared to VOS but has been gaining interest.
-We evaluate our method on the dense-video task of the UVO v1.01.01.01.0 [34] dataset.
-[…]
-The Unidentified Video Objects (UVO) dataset is designed to recognize and segment all objects regardless of the categories, even those unseen during training, thereby focusing on VIS in the open world.
-Each video in UVO features on average 12.3 object annotations, a considerable increase from previous datasets having only 2 or 3 objects per video on average.
-UVO sources its videos from the Kinetics-400 [14] dataset and contains three different splits: FrameSet, VideoSparseSet, and VideoDenseSet.
-The VideoDenseSet consists of 3-second clips annotated densely at 30fps and tracked over time.
-The primary goal of VideoDenseSet is to study video open-world segmentation.
-Objects identifiable under COCO categories carry their respective COCO labels, while ambiguous objects or those outside the COCO taxonomy are labeled as “other”.
-This meticulous and exhaustive annotation structure makes the VideoDenseSet ideal for research areas that require an understanding of videos in a dense and comprehensive manner, such as robotics, autonomous driving, and augmented-reality applications.
-Metrics.
-We evaluate our method using standard evaluation metrics in image instance segmentation, adapted for video instance segmentation [38].
-These include Average Precision (AP) and Average Recall (AR) IoU-based metrics.
-Given that each instance in a video comprises a sequence of masks, unlike image instance segmentation, IoU computation is carried out not only in the spatial dimensions but also in the temporal dimension.
-This implies that the sum of intersections at every single frame is divided by the sum of unions at every single frame.
-These metrics are generally computed on a per-category basis and subsequently averaged across all categories.
-However, we work with the class-agnostic version of UVO.
+DAVIS $2017$ is also used in our interactive point-based video segmentation study.
+We additionally devise a VOS dataset from BDD100K [50].
+For VIS, We evaluate our method on the class-agnostic dense video instance segmentation task of the UVO v1.01.01.01.0 [40] dataset.
+UVO v1.01.01.01.0 is a VIS dataset aiming for open-world segmentation, where objects of any category, including those unseen in training, are identified and segmented.
 4.2 Implementation Details
 Training Data.
-For our experiments, we use pre-trained checkpoints provided by the respective authors for both PIPS [11] and SAM.
-PIPS is trained exclusively on a synthetic dataset, FlyingThings++ [11], derived from the FlyingThings [24] optical flow dataset.
-This dataset includes multi-frame amodal trajectories with synthetic occlusions caused by moving objects.
-SAM, on the other hand, has been trained on the large-scale SA-1B dataset, the largest image segmentation dataset to date, with over 1 billion masks on 11M licensed and privacy-respecting images.
-It is noteworthy that neither of these datasets includes video segmentation data, and they do not overlap with any of our evaluation data.
-This effectively positions our model in a zero-shot video segmentation setting.
-Model Variations.
-Our experiments led to two optimal model hyperparameters, distinguished as SAM-PT (without reinitialization) and SAM-PT-reinit (with reinitialization).
-These configurations were derived from our ablation study in Sec.
-4.3.
-However, we found that using iterative refinement negatively impacted both SAM-PT and SAM-PT-reinit on the MOSE dataset, and likewise hindered SAM-PT-reinit on the YouTube-VOS dataset.
-Consequently, iterative refinement was deactivated for these specific datasets.
-For DAVIS, we additionally report results for replacing SAM with HQ-SAM [15] and denote the model variants as HQ-SAM-PT and HQ-SAM-PT-reinit.
-The HQ-SAM variants use 3 iterative refinement iterations instead of 12 iterations.
-VOS Evaluation.
+For our experiments, we use pre-trained checkpoints provided by the respective authors for the point trackers (PIPS [32], CoTracker [19], etc.) and SAM.
+PIPS and CoTracker have been trained exclusively on synthetic data, PIPS on FlyingThings++ [16] and CoTracker on TAP-Vid-Kubric [14].
+SAM has been trained on the large-scale SA-1B dataset, the largest image segmentation dataset to date.
+HQ-SAM is further trained on the HQ-Seg-44k [20].
+Noteworthy, none of these datasets contain video segmentation data, nor do they intersect with any datasets we use for evaluation, situating our model within a zero-shot setting.
+Interactive Point-Based Video Segmentation.
+To assess the interactive capabilities of SAM-PT, we simulate user refinement of video segmentation results through point additions and removals.
+We compare three methods: a non-tracking approach using SAM alone, an online method making one pass through the video with a target IoU quality, and an offline method employing multiple passes that progressively aim at higher IoU quality.
+These methods are detailed in the Supplementary Material, which includes pseudocode for each simulation variant.
+4.3 Ablation Study
+Our ablation experiments on the DAVIS $2017$ validation subset assessed different aspects of SAM-PT’s design.
+Despite the valuable insights, we acknowledge the dataset’s limited scope in representing diverse and complex segmentation challenges, such as occlusions and varying environmental conditions, may constrain the generalizability of our ablation’s findings.
+Future investigations could benefit from a more varied validation set, potentially sourced from the YouTube-VOS $2018$ training dataset, to enhance robustness.
+Our findings, detailed in Tab.
+2, underscore SAM-PT’s adaptability across leading long-term point trackers – PIPS [16], TAPIR [15], and notably CoTracker [19], which excelled due to its precise point tracking and reliable occlusion predictions.
+In contrast, PIPS++ [53] lagged despite being a more recent iteration of PIPS[16], due to the lack of occlusion prediction which is important for the effective use of point tracking for segmentation.
+TapNet [14] struggled due to less effective temporal consistency and high-resolution inputs.
+Traditional methods such as SuperGlue [33] and RAFT [35], which, although proficient in their respective domains, either struggle with the dynamic and deformable aspects of video scenes or cannot handle occlusion, highlighting the specialized efficacy of long-term trackers in the video segmentation landscape.
+Table 2: We report the mean performance and standard deviation across eight runs on the validation subset of DAVIS $2017$ to study the impact of different point trackers.
+Point Tracker DAVIS 2017 Validation [30]
+[…]
+SuperGlue [33] $28.4$±$\pm 3.1$ $\pm 2.4$ $32.0$±$\pm 3.8$
+TapNet [14] $60.9$±$\pm 0.2$ $58.2$±$\pm 0.3$ $63.5$±$\pm 0.2$
+RAFT [35] $63.0$±$\pm 0.6$ $\pm 0.6$ $65.4$±$\pm 0.5$
+PIPS++ [53] $73.2$±$\pm 0.5$ $69.9$±$\pm 0.5$ $76.6$±$\pm 0.5$
+PIPS [16] $76.3$±$\pm 0.6$ $73.6$±$\pm 0.6$ $78.9$±$\pm 0.6$
+TAPIR [15] $76.7$±$\pm 0.3$ $73.8$±$\pm 0.4$ $79.7$±$\pm 0.3$
+CoTracker [19] $\mathbf{77.6}$±$\pm 0.7$ $\mathbf{74.8}$±$\pm 0.7$ $\mathbf{80.4}$±$\pm 0.7$
+In Tab.
+3, we tested SAM-PT with various settings using PIPS as the tracker.
+We found that using eight positive points per object instead of just one improved our scores significantly by $33.4$ points because one point often wasn’t enough for unambiguously prompting SAM.
+Selecting points with K-Medoids was slightly better than random and matched Shi-Tomasi, giving a boost of $1.8$ points.
+Incorporating negative points besides positive points helped when trackers made mistakes, such as losing track of an object, by improving scores by another $1.8$ points.
+Adding iterative refinement smoothed out mask quality and fixed some errors, adding another $2.2$ points to our performance.
+We tried filtering out unreliable points with patch similarity, but this did not work well as it ended up removing too many points.
+Finally, although reinitializing points did not help significantly in the initial tests, it did show benefits on other datasets such as MOSE and UVO, helping us recover from tracker errors by discarding incorrect and adding fresh points as well as detecting that the object has disappeared and the tracking should be halted.
+Table 3: Ablation study on the validation subset of DAVIS $2017$ on the impact of different SAM-PT configurations using PIPS [16] as the point tracker.
+PSM: point selection method.
+PP: positive points per mask.
+NP: negative points per mask.
+IRI: iterative refinement iterations.
+PS: point similarity filtering.
+RV: reinitalization variant.
+SAM-PT Configuration (using PIPS) DAVIS [30]
+PSM PP NP IRI PS RV $\mathcal{J\&F}$ Gain
+Random $1$ 00 00 $37.1$±$\pm 21.7$
+Random $8$ 00 00 $70.5$±$\pm 1.4$ $+33.4$
+K-Medoids $8$ 00 00 $72.3$±$\pm 1.2$ $+1.8$
+Shi-Tomasi $8$ 00 00 $72.0$±$\pm 0.3$
+Mixed $8$ 00 00 $70.6$±$\pm 0.8$
+K-Medoids $8$ $1$ 00 $74.1$±$\pm 0.7$ $+1.8$
+K-Medoids $8$ $1$ $12$ $76.3$±$\pm 0.6$ $+2.2$
+K-Medoids $8$ $1$ $12$ $72.7$±$\pm 2.0$ none
+K-Medoids $8$ $72$ $12$ A $76.8$±$\pm 0.7$ $+0.5$
+K-Medoids $8$ $1$ $12$ B $76.1$±$\pm 0.4$
+K-Medoids $8$ $1$ 00 C $75.5$±$\pm 0.7$
+K-Medoids $8$ $1$ $12$ D $76.4$±$\pm 0.3$
+Extended ablation experiments and discussions can be found in the Supplementary Material, including complete ablation results for PIPS and CoTracker, the choice of the SAM backbone, and SAM’s lightweight variants that reveal trade-offs between performance and inference speed.
+4.4 Video Object Segmentation
+Performance Overview.
+Our SAM-PT method, utilizing HQ-SAM and CoTracker, sets a new standard in zero-shot video object segmentation on the DAVIS $2017$ dataset with a mean $\mathcal{J\&F}$ score of $79.4$, outperforming SegGPT’s $75.6$, DINO’s $71.4$, and Painter’s $34.6$ as shown in Tab.
+4.
+On the easier DAVIS $2016$ validation set, our method achieves $84.3$, surpassing SegGPT’s $82.3$, showcasing the strength of our approach even in less complex scenarios, as detailed in the Supplementary Material.
+Table 4: Quantitative results on the DAVIS $2017$ [30] validation set for semi-supervised VOS.
+Performance is reported for different methods, including our SAM-PT and HQ-SAM-PT, with and without the reinitialization strategy (Reinit) and using different point trackers.
+Our method outperforms other zero-shot methods.
+Method Tracker Reinit DAVIS $\mathbf{2017}$ Validation [30]
+[…]
+(a) trained on video segmentation data
+MiVOS [8] - - $84.5$ $81.7$ $87.4$
+DeAOT [48] - - $86.2$ $83.1$ $89.2$
+DEVA [10] - - $87.6$ $84.2$ $91.0$
+XMem [7] - - $87.7$ $84.0$ $91.4$
+(b) not trained on video segmentation data (zero-shot)
+Painter [41] - - $34.6$ $28.5$ $40.8$
+DINO [4] - - $71.4$ $67.9$ $74.9$
+SegGPT [42] - - 75.675.6{75.6}75.6 72.572.5{72.5}72.5 78.678.6{78.6}78.6
+SAM-PT (ours) PIPS [16] $76.3$±$\pm 0.6$ $73.6$±$\pm 0.6$ 78.978.9{78.9}78.9±$\pm 0.6$
+PIPS [16] 76.676.6{76.6}76.6±$\pm 0.7$ 74.474.4{74.4}74.4±$\pm 0.8$ 78.978.9{78.9}78.9±$\pm 0.6$
+CoTracker [19] 77.677.6{77.6}77.6±$\pm 0.7$ 74.874.8{74.8}74.8±$\pm 0.7$ 80.480.4{80.4}80.4±$\pm 0.7$
+CoTracker [19] $77.4$±$\pm 1.0$ $74.5$±$\pm 1.0$ $80.3$±$\pm 1.1$
+HQ-SAM-PT (ours) PIPS [16] $77.2$±$\pm 0.5$ $74.7$±$\pm 0.5$ $79.8$±$\pm 0.4$
+PIPS [16] $77.0$±$\pm 0.7$ $74.8$±$\pm 0.8$ $79.2$±$\pm 0.6$
+CoTracker [19] 79.4±$\pm 0.6$ 76.5±$\pm 0.6$ 82.3±$\pm 0.5$
+CoTracker [19] $77.7$±$\pm 0.8$ $74.6$±$\pm 0.9$ $80.8$±$\pm 0.7$
+For the YouTube-VOS $2018$ validation set, we achieve the highest performance among zero-shot methods with $76.2$ against SegGPT’s $74.7$ and Painter’s $24.1$, indicating robust generalizability across various video segmentation benchmarks (Tab.
+5).
+Table 5: Quantitative results in semi-supervised VOS on the validation subset of YouTube-VOS $2018$.
+Metrics are reported separately for “seen” and “unseen” classes, with $\mathcal{G}$ being the overall average score over the metrics.
+Method Tracker Reinit YouTube-VOS $\mathbf{2018}$ Validation [44]
+$\mathcal{G}$ $\mathcal{J}_{s}$ $\mathcal{F}_{s}$ $\mathcal{J}_{u}$ $\mathcal{F}_{u}$
+(a) trained on video segmentation data
+XMem [7] - - $86.1$ $85.1$ $89.8$ $80.3$ $89.2$
+DeAOT [48] - - $86.2$ $85.6$ $90.6$ $80.0$ $88.4$
+(b) not trained on video segmentation data (zero-shot)
+Painter [41] - - $24.1$ $27.6$ $35.8$ $14.3$ $18.7$
+SegGPT [42] - - 74.774.7{74.7}74.7 75.175.1{75.1}75.1 $\mathbf{80.2}$ 67.467.4{67.4}67.4 75.975.9{75.9}75.9
+SAM-PT (ours) PIPS [16] $67.0{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.3}}$ $68.6{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.2}}$ $71.2{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.1}}$ $61.0{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.5}}$ $67.4{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.4}}$
+PIPS [16] ${67.5}{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.2}}$ ${69.0}{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.4}}$ ${69.9}{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.3}}$ ${63.2}{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.4}}$ ${67.8}{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.5}}$
+CoTracker [19] $74.0{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.3}}$ $73.3{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.2}}$ $76.0{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.2}}$ $70.0{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.4}}$ $76.7{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.4}}$
+CoTracker [19] $71.5{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.4}}$ $71.0{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.3}}$ $72.8{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.3}}$ $68.3{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.7}}$ $73.9{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.7}}$
+HQ-SAM-PT (ours) CoTracker [19] $\mathbf{76.2}{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.1}}$ $\mathbf{75.3}{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.1}}$ $78.4{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.2}}$ $\mathbf{72.1}{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.2}}$ $\mathbf{79.0}{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.2}}$
+In the semi-supervised VOS on BDD100K’s validation set, our method outperforms SegGPT for non-transient objects but also surpasses the fully-supervised XMem across nearly all object visibility durations.
+The detailed breakdown is provided in Tab.
+6.
+On the MOSE $2023$ validation set, our performance remains competitive with SegGPT, with exact figures available in the Supplementary Material.
+(a) Successful segmentation cases for SAM-PT using 8 positive and 1 negative point.
+(b) Successful segmentation cases for SAM-PT with 8 positive and 72 negative points and reinitialization enabled.
+(c) Failure cases for SAM-PT where challenges such as occlusions and thin object structures lead to tracking errors.
+Figure 5: Visualization of SAM-PT on DAVIS 2017 [30].
+The method shows its capability to segment and track objects using the initial masks from the first frame, with circles denoting positive points and crosses negative points.
+Red symbols indicate occlusion prediction.
+Figure 6: Successful segmentation using SAM-PT on short clips from “Avatar: The Last Airbender”.
+Although our method has never seen data from Avatar, an anime-influenced animated television series, it segments and tracks various objects in short clips.
+Qualitative Analysis.
+Our visualizations in Fig.
+4(a) and Fig.
+4(b) demonstrate successful segmentation on the DAVIS $2017$ dataset and underscore our method’s ability to perform zero-shot video segmentation on unseen content, such as clips from the anime-influenced series “Avatar: The Last Airbender” in Fig.
+6.
+These examples highlight the versatility and adaptability of SAM-PT.
+Limitations and Challenges.
+Our method excels in zero-shot video object segmentation but faces challenges with point tracker reliability in complex scenarios, such as occlusions and fast-moving objects, as shown in Fig.
+4(c).
+While point reinitialization and negative point strategies offer some improvement, interactive use significantly bridges the performance gap with trained methods.
+This interactive potential is successfully demonstrated in Sec.
+4.6, where user intervention enhances segmentation accuracy.
+4.5 Video Instance Segmentation
+Given the same mask proposals, SAM-PT outperforms TAM [46] significantly, as shown in Tab.
+7, even though SAM-PT was not trained on any video segmentation data.
+TAM is a concurrent approach combining SAM and XMem [7], where XMem was pre-trained on BL30K [9] and trained on DAVIS and YouTube-VOS, but not on UVO.
+On the other hand, SAM-PT combines SAM with the PIPS or CoTracker point tracking method, both of which have not been trained on video segmentation tasks.
+4.6 Interactive Point-Based Video Segmentation
+Building upon our method’s strengths observed in standard benchmarks, this study evaluates the responsiveness of SAM-PT to human input, aiming to understand its practical utility for interactive video annotation tasks.
+We benchmarked SAM-PT’s interactive performance against a baseline SAM approach that does not utilize point tracking.
+The results, visualized in Fig.
+7, show that SAM-PT significantly outperforms the baseline, particularly when employing the offline checkpoint strategy.
+Although this offline method initially starts at a lower performance due to setup costs, it quickly exceeds the online method’s performance by benefiting from a cumulative optimization approach.
+Figure 7: Interactive segmentation performance on the DAVIS 2017 [30] validation set.
+SAM-PT’s offline strategy notably outperforms the baseline SAM, demonstrating efficient video annotation with minimal user intervention.
+These results suggest that SAM-PT substantially reduces the effort required for high-quality video annotation, bringing its performance closer to fully-supervised methods and highlighting its practical utility.
+Table 6: This table shows the semi-supervised VOS performance on BDD100K’s validation set, comparing our HQ-SAM-PT method using CoTracker as the point tracker (without reinitialization) against the zero-shot SegGPT and the fully-supervised XMem.
+Performance metrics include the average $\mathcal{J\&F}$ measure for object visibility durations categorized as short (1-5 frames), medium (6-30 frames), and long (31+ frames).
+Our approach demonstrates superior results over SegGPT for non-transient objects and over XMem across all visibility durations except for long-term object tracking.
+BDD100K VOS Validation [50]
+[…]
+(a) trained on video segmentation data, but not on BDD100K
+XMem [7] $76.6$ $74.5$ $78.7$ $79.3$ $78.6$ $\mathbf{63.7}$
+(b) not trained on video segmentation data (zero-shot)
+SegGPT [42] $\mathbf{81.5}$ $\mathbf{81.2}$ $\mathbf{81.8}$ $\mathbf{96.1}$ $\underline{78.6}$ $52.0$
+HQ-SAM-PT (ours) $\underline{81.0}$ $\underline{80.1}$ $\mathbf{81.8}$ $\underline{91.8}$ $\mathbf{79.9}$ $\underline{55.8}$
+[…]
+Method Tracker Reinit $\mathbf{AR100}$ $\mathbf{ARs}$ $\mathbf{ARm}$ $\mathbf{ARl}$ $\mathbf{AP}$
+(a) trained on video segmentation data, including UVO’s training subset
+Mask2Former VIS [51] - - $35.4$- -- -- $27.3$
+ROVIS [51] - - $41.2$- -- -- $32.7$
+(b) trained on video segmentation data
+TAM [46] - - $24.1$ $21.1$ $32.9$ $31.1$ $1.7$
+(c) not trained on video segmentation data (zero-shot)
+SAM-PT (ours) PIPS [16] $28.8$ $23.3$ $40.8$ $48.3$ $\mathbf{6.7}$
+PIPS [16] $\mathbf{30.8}$ $\mathbf{25.1}$ $\mathbf{44.1}$ $\mathbf{49.2}$ $6.5$
+CoTracker [19] $29.5$ $25.3$ $39.0$ $44.1$ $5.8$
+CoTracker [19] $29.8$ $25.1$ $40.6$ $45.6$ $6.2$
+5 Conclusion
+SAM-PT introduces a point-centric approach for interactive video segmentation by combining the generalization capabilities of the Segment Anything Model (SAM) and long-term point tracking.
+Our work fills in the gap that the point-centric approach is scantly explored in the literature.
+In our experiments, SAM-PT achieves strong performance across video segmentation tasks including semi-supervised, open-world, and fully interactive video segmentation.
+While our method has limitations such as difficulty handling occlusions, small objects, motion blur, and inconsistencies in mask predictions, it contributes a new perspective to video object segmentation research.
+Supplementary Material
+In this supplementary, we first report experimental results on additional datasets and subsets (Appendices A, B and C).
+Then we extend and detail our ablation and report more qualitative results (Appendices D, E, F, G and H).
+Lastly, we detail on our evaluation protocols (Appendices I, J, K and L).
+Appendix A MOSE 2023
+MOSE $2023$ [13] is a recently introduced dataset that focuses on multi-object segmentation and tracking in complex scenes, replete with challenges such as occlusions, transient visibility of objects, extensive occlusion, etc.
+Our results on the validation subset suggest that SAM-PT achieves performance competitive with SegGPT [42], as shown in Tab.
+8.
+Table 8: Quantitative results on the MOSE $2023$ validation set for semi-supervised VOS.
+Our method achieves performance comparable to the state-of-the-art zero-shot learning method.
+Note that SegGPT and SAM-PT adopt completely different training data.
+Method Tracker Reinit MOSE $\mathbf{2023}$ Validation [13]
+[…]
+(a) trained on video segmentation data
+RDE [23] - - $48.8$ $44.6$ $52.9$
+XMem [7] - - $57.6$ $53.3$ $62.0$
+DeAOT [48] - - $59.4$ $55.1$ $63.8$
+DEVA [10] - - $66.5$ $62.3$ $70.8$
+(b) not trained on video segmentation data (zero-shot)
+Painter [41] - - $14.5$ $10.4$ $18.5$
+SegGPT [42] - - $\mathbf{45.1}$ $\mathbf{42.2}$ $\mathbf{48.0}$
+SAM-PT (ours) PIPS [16] $38.5{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.2}}$ $34.9{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.3}}$ $42.1{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.2}}$
+PIPS [16] ${41.0}{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.5}}$ ${38.5}{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.5}}$ ${43.5}{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.5}}$
+CoTracker [19] $41.8{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.2}}$ $38.3{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.2}}$ $45.2{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.3}}$
+CoTracker [19] $40.1{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.5}}$ $36.0{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.5}}$ $44.1{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.4}}$
+TAPIR [15] $\underline{42.9}{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.2}}$ $38.3{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.2}}$ $\underline{47.6}{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.1}}$
+HQ-SAM-PT (ours) CoTracker [19] $42.4$±${\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.3}}$ $\underline{39.0}$±${\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.3}}$ $45.8$±${\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.3}}$
+TAPIR [15] $42.1{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.1}}$ $37.6{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.1}}$ $46.7{\color[rgb]{.5,.5,.5}\scriptsize{\pm 0.1}}$
+Appendix B DAVIS 2016
+DAVIS $2016$ offers a single-object VOS benchmark across $20$ diverse sequences.
+We report results on the DAVIS $2016$ validation subset in Tab.
+9, in which our method achieves $84.3$ points, surpassing SegGPT’s $82.3$ points.
+Table 9: Quantitative results on the DAVIS $2016$ validation set for semi-supervised VOS. Our method achieves higher performance compared to SegGPT, both of which are zero-shot methods.
+DAVIS 2016 Validation [30]
+[…]
+SegGPT [42] - - $82.3$ $81.8$ $82.8$
+SAM-PT PIPS [16] 83.183.1{83.1}83.1±$\pm 1.5$ 83.083.0{83.0}83.0±$\pm 0.8$ 83.083.0{83.0}83.0±$\pm 1.1$
+PIPS [16] $80.2$±$\pm 0.6$ $80.3$±$\pm 0.6$ $80.0$±$\pm 0.6$
+CoTracker [19] $83.1\pm 0.6$ $83.2\pm 0.7$ $82.9\pm 0.6$
+CoTracker [19] $82.6\pm 0.8$ $83.0\pm 1.0$ $82.2\pm 0.9$
+HQ-SAM-PT CoTracker [19] $\mathbf{84.3}\pm 0.9$ $\mathbf{84.9}\pm 1.0$ $\mathbf{83.7}\pm 0.9$
+Appendix C DAVIS 2017 Test-dev Subset
+DAVIS $2017$ is a multi-object extension of its $2016$ version.
+The video scenarios within this dataset are small but diverse.
+In addition to the results on the validation subset in the main manuscript, we report the performance of SAM-PT on the DAVIS $2017$ test-dev subset in Tab.
+10.
+Table 10: Quantitative results on the DAVIS 2017 test-dev subset for semi-supervised VOS.
+DAVIS 2017 Test-dev [30]
+[…]
+SAM-PT PIPS [16] 62.762.7{62.7}62.7±$\pm 0.5$ 59.459.4{59.4}59.4±$\pm 0.6$ 66.166.1{66.1}66.1±$\pm 0.4$
+PIPS [16] $61.5$±$\pm 1.1$ $59.3$±$\pm 1.0$ $63.8$±$\pm 1.2$
+CoTracker [19] $65.7$±$\pm 0.7$ $62.8$±$\pm 0.7$ $68.5$±$\pm 0.7$
+CoTracker [19] $62.0$±$\pm 1.2$ $58.8$±$\pm 1.2$ $65.1$±$\pm 1.1$
+TAPIR [15] $\mathbf{69.0}$ $\mathbf{66.0}$ $\mathbf{72.0}$
+TAPIR [15] $64.1$ $61.3$ $66.9$
+HQ-SAM-PT CoTracker [19] $64.8$±$\pm 0.5$ $61.9$±$\pm 0.5$ $67.7$±$\pm 0.5$
+Appendix D Different SAM Backbones
+The SAM model’s backbone plays an important role in determining its performance and inference speed.
+In this experiment, we evaluated SAM with different ViT backbones: ViT-Huge (used throughout the work), ViT-Large, and ViT-Base.
+The results, as measured on the validation subset of DAVIS $2017$ are shown in Tab.
+11.
+Replacing ViT-Huge with ViT-Large results in only a non-significant loss in performance, SAM-PT’s overall performance nevertheless remains non-real-time.
+Note that the ViT-Huge number in this experiment is slightly different from the one presented in the main manuscript due to the use of different seed values.
+Table 11: Performance of SAM-PT with different backbones and their inference speed in semi-supervised VOS on the validation subset of DAVIS $2017$, when using PIPS as the point tracker.
+SAM Backbone $\mathcal{J\&F}$ FPS
+ViT-Huge $\mathbf{76.7}\pm 0.6$ $1.4$
+ViT-Large $76.4\pm 0.6$ $1.8$
+ViT-Base $72.2\pm 0.5$ $\mathbf{2.6}$
+Appendix E Different SAM Variants
+To cater to scenarios where inference speed is crucial, we explored lightweight variants of SAM: Light HQ-SAM [20] and MobileSAM [52].
+The performance and speed trade-offs for these variants are summarized in table Tab.
+12.
+Using the HQ-SAM [20] variant of SAM results in the highest performance of $77.64$ points, whereas MobileSAM has the highest inference speed of $5.5$ FPS.
+Using the lightweight variants doesn’t achieve real-time performance as the bottleneck of the pipeline moves to the point tracker.
+Table 12: Performance of lightweight SAM variants and their inference speed in semi-supervised VOS on the validation subset of DAVIS $2017$ when using PIPS [16] as the point tracker.
+SAM Variant Backbone $\mathcal{J\&F}$ FPS
+HQ-SAM [20] ViT-Huge $\mathbf{77.64}$ $1.3$
+SAM [21] ViT-Huge $76.65$ $1.4$
+Light HQ-SAM [20] ViT-Tiny $71.30$ $4.8$
+MobileSAM [52] ViT-Tiny $71.07$ $\mathbf{5.5}$
+Appendix F Detailed Ablation Results
+We report detailed experimental results of our ablation studies for configurations using PIPS [16] as the point tracker in Tab. 14 and using CoTracker [19] in Tab. 13.
+Table 13: Ablation study on the validation subset of DAVIS $2017$ on the impact of different SAM-PT configurations using CoTracker [19] as the point tracker.
+PSM: point selection method.
+PP: positive points per mask.
+NP: negative points per mask.
+IRI: iterative refinement iterations.
+PS: point similarity filtering.
+SAM-PT Config. (using CoTracker) DAVIS 2017 Validation [30]
+[…]
+K-Medoids 16 1 12 $\mathbf{77.6}$±$\pm 0.7$ $\mathbf{74.8}$±$\pm 0.7$ $\mathbf{80.4}$±$\pm 0.7$
+Shi-Tomasi 16 1 12 $74.3$±$\pm 0.3$ $72.1$±$\pm 0.3$ $76.5$±$\pm 0.3$
+Random 16 1 12 $76.4$±$\pm 1.1$ $73.3$±$\pm 1.1$ $79.4$±$\pm 1.0$
+Mixed 16 1 12 $76.4$±$\pm 0.6$ $73.7$±$\pm 0.5$ $79.2$±$\pm 0.6$
+K-Medoids 1 1 12 $39.0$±$\pm 0.9$ $36.1$±$\pm 0.9$ $42.0$±$\pm 1.0$
+K-Medoids 16 0 12 $76.8$±$\pm 0.7$ $74.1$±$\pm 0.7$ $79.6$±$\pm 0.7$
+K-Medoids 16 1 0 $75.3$±$\pm 0.6$ $73.2$±$\pm 0.5$ $77.3$±$\pm 0.6$
+K-Medoids 16 1 1 $76.6$±$\pm 0.6$ $74.3$±$\pm 0.6$ $78.9$±$\pm 0.6$
+K-Medoids 16 1 100 $77.5$±$\pm 0.7$ $74.7$±$\pm 0.7$ $80.3$±$\pm 0.7$
+K-Medoids 16 1 12 $73.8$±$\pm 0.8$ $71.1$±$\pm 0.8$ $76.5$±$\pm 0.8$
+Table 14: Detailed ablation study on the validation subset of DAVIS $2017$ on the impact of different SAM-PT configurations using PIPS [16] as the point tracker.
+PSM: point selection method.
+PP: positive points per mask.
+NP: negative points per mask.
+IRI: iterative refinement iterations.
+PS: point similarity filtering.
+RV: reinitialization variant.
+SAM-PT Configuration (using PIPS) DAVIS $\mathbf{2017}$ Validation [30]
+[…]
+(a) point selection method and positive points per mask
+Random $1$ 00 00 $37.1$±$\pm 21.7$ $34.3$±$\pm 22.0$ $40.0$±$\pm 21.5$
+Random $8$ 00 00 $70.5$±$\pm 1.4$ $68.5$±$\pm 1.4$ $72.6$±$\pm 1.5$
+Random $16$ 00 00 $70.0$±$\pm 1.1$ $68.2$±$\pm 1.0$ $71.8$±$\pm 1.2$
+Random $72$ 00 00 $62.6$±$\pm 0.4$ $62.3$±$\pm 0.3$ $62.8$±$\pm 0.5$
+Shi-Tomasi $1$ 00 00 $20.3$±$\pm 0.1$ $18.3$±$\pm 0.1$ $22.3$±$\pm 0.2$
+Shi-Tomasi $8$ 00 00 $72.0$±$\pm 0.3$ $70.3$±$\pm 0.3$ $73.7$±$\pm 0.4$
+Shi-Tomasi $16$ 00 00 $66.6$±$\pm 0.4$ $65.7$±$\pm 0.5$ $67.6$±$\pm 0.4$
+Shi-Tomasi $72$ 00 00 $54.4$±$\pm 0.3$ $54.8$±$\pm 0.2$ $54.0$±$\pm 0.4$
+K-Medoids $1$ 00 00 $32.2$±$\pm 0.7$ $30.4$±$\pm 0.8$ $34.0$±$\pm 0.7$
+K-Medoids $8$ 00 00 $\mathbf{72.3}$±$\pm 1.2$ $\mathbf{70.4}$±$\pm 1.3$ $\mathbf{74.3}$±$\pm 1.1$
+K-Medoids $16$ 00 00 $71.4$±$\pm 0.2$ $69.8$±$\pm 0.3$ $73.1$±$\pm 0.2$
+K-Medoids $72$ 00 00 $58.0$±$\pm 0.2$ $57.3$±$\pm 0.2$ $58.7$±$\pm 0.3$
+Mixed $1$ 00 00 $29.9$±$\pm 0.9$ $26.6$±$\pm 0.8$ $33.2$±$\pm 1.4$
+Mixed $8$ 00 00 $70.6$±$\pm 0.8$ $68.6$±$\pm 0.8$ $72.5$±$\pm 0.8$
+Mixed $16$ 00 00 $\pm 0.7$ $68.2$±$\pm 0.6$ $71.9$±$\pm 0.7$
+Mixed $72$ 00 00 $62.8$±$\pm 0.5$ $62.4$±$\pm 0.5$ $63.2$±$\pm 0.6$
+(b) negative points per mask
+K-Medoids $8$ 00 00 $72.3$±$\pm 1.2$ $70.4$±$\pm 1.3$ $74.3$±$\pm 1.1$
+K-Medoids $8$ $1$ 00 $\mathbf{74.1}$±$\pm 0.7$ $\mathbf{72.1}$±$\pm 0.6$ $\mathbf{76.1}$±$\pm 0.7$ $\mathbf{+1.8}$
+K-Medoids $8$ $8$ 00 $74.0$±$\pm 0.8$ $71.9$±$\pm 0.8$ $76.0$±$\pm 0.9$
+K-Medoids $8$ $16$ 00 $73.4$±$\pm 0.6$ $71.4$±$\pm 0.6$ $75.3$±$\pm 0.6$
+K-Medoids $8$ $72$ 00 $72.2$±$\pm 0.4$ $70.3$±$\pm 0.4$ $74.0$±$\pm 0.4$
+(c) iterative refinement iterations
+K-Medoids $8$ $1$ 00 $74.1$±$\pm 0.7$ $72.1$±$\pm 0.6$ $76.1$±$\pm 0.7$
+K-Medoids $8$ $1$ $1$ $75.7$±$\pm 0.7$ $73.4$±$\pm 0.7$ $78.1$±$\pm 0.6$
+K-Medoids $8$ $1$ $3$ $76.0$±$\pm 0.6$ $73.4$±$\pm 0.7$ $78.6$±$\pm 0.7$
+K-Medoids $8$ $1$ $12$ $\mathbf{76.3}$±$\pm 0.6$ $\mathbf{73.6}$±$\pm 0.6$ $\mathbf{78.9}$±$\pm 0.6$ $\mathbf{+2.2}$
+(d) patch similarity filtering
+K-Medoids $8$ $1$ $12$ $\mathbf{76.3}$±$\pm 0.6$ $\mathbf{73.6}$±$\pm 0.6$ $\mathbf{78.9}$±$\pm 0.6$ none
+K-Medoids $8$ $1$ $12$ $0.002$ $72.7$±$\pm 2.0$ $70.2$±$\pm 1.8$ $75.2$±$\pm 2.1$
+K-Medoids $8$ $1$ $12$ $0.01$ $70.7$±$\pm 2.0$ $68.3$±$\pm 1.8$ $73.2$±$\pm 2.1$
+(e) point reinitialization
+K-Medoids $8$ $1$ 00 A $75.7$±$\pm 0.7$ $73.7$±$\pm 0.6$ $77.7$±$\pm 0.8$
+K-Medoids $8$ $1$ 00 B $75.8$±$\pm 0.6$ $73.5$±$\pm 0.9$ $78.1$±$\pm 0.3$
+K-Medoids $8$ $1$ 00 C $75.5$±$\pm 0.7$ $73.2$±$\pm 0.8$ $77.8$±$\pm 0.7$
+K-Medoids $8$ $1$ 00 D $75.4$±$\pm 0.2$ $73.3$±$\pm 0.2$ $77.5$±$\pm 0.3$
+K-Medoids $8$ $1$ $12$ A $76.6$±$\pm 0.8$ $74.0$±$\pm 0.8$ $\mathbf{79.1}$±$\pm 0.8$
+K-Medoids $8$ $1$ $12$ B $76.1$±$\pm 0.4$ $73.5$±$\pm 0.5$ $78.6$±$\pm 0.3$
+K-Medoids $8$ $1$ $12$ C $75.4$±$\pm 0.6$ $72.8$±$\pm 0.7$ $78.0$±$\pm 0.5$
+K-Medoids $8$ $1$ $12$ D $76.4$±$\pm 0.3$ $74.0$±$\pm 0.4$ $78.8$±$\pm 0.3$
+K-Medoids $8$ $1$ $12$ $76.3$±$\pm 0.6$ $73.6$±$\pm 0.6$ $78.9$±$\pm 0.6$
+K-Medoids $8$ $72$ 00 A $74.9$±$\pm 0.9$ $73.2$±$\pm 0.8$ $76.6$±$\pm 1.0$
+K-Medoids $8$ $72$ 00 B $76.0$±$\pm 1.1$ $73.9$±$\pm 1.1$ $78.1$±$\pm 1.1$
+K-Medoids $8$ $72$ 00 C $75.1$±$\pm 0.6$ $72.9$±$\pm 0.5$ $77.2$±$\pm 0.7$
+K-Medoids $8$ $72$ 00 D $75.6$±$\pm 1.5$ $73.8$±$\pm 1.5$ $77.3$±$\pm 1.6$
+K-Medoids $8$ $72$ $12$ A $\mathbf{76.8}$±$\pm 0.7$ $\mathbf{74.5}$±$\pm 0.8$ 79.079.0{79.0}79.0±$\pm 0.6$ $\mathbf{+0.5}$
+K-Medoids $8$ $72$ $12$ B $74.8$±$\pm 0.8$ $72.1$±$\pm 0.9$ $77.6$±$\pm 0.7$
+K-Medoids $8$ $72$ $12$ C $75.0$±$\pm 0.4$ $72.1$±$\pm 0.4$ $77.8$±$\pm 0.5$
+K-Medoids $8$ $72$ $12$ D $75.2$±$\pm 1.1$ $72.7$±$\pm 1.1$ $77.6$±$\pm 1.1$
+Appendix G Point Tracking Reinitialization
+In our method, we introduce an optional reinitialization strategy.
+Here, the point tracker begins anew every $h$ frames, where $h$ represents a pre-set tracking horizon (e.g., $8$ frames), or is dynamically determined based on SAM’s mask predictions for each timestep within the horizon (e.g., using most-similar-mask-area heuristics).
+Upon reaching this horizon, the query points given to the tracker are reinitialized according to the mask prediction SAM outputted at the horizon frame.
+While this method may increase the computational load, it shows performance improvement when using the PIPS [16] point tracker in demanding video sequences, such as those in the MOSE dataset.
+However, our studies also suggested that the proposed reinitialization strategies hurt performance when using CoTracker as the point tracker.
+We primarily designed the reinitialization variants to address some of the failure cases of PIPS such as the common case of points being wrongly predicted to be off the target object, being predicted on the background instead, but this does not work as well for other point trackers such as CoTracker that are more robust to such failures.
+We explored four reinitialization strategies, each varying in how they compute the value of $h$:
+  1. (A)
+    Reinit-on-Horizon-and-Sync-Masks: This straightforward variant reinitializes points after a fixed number of frames (e.g., every 8 frames).
+However, it may stumble if the mask is absent at the reinitialization timestep.
+  2. (B)
+    Reinit-at-Median-of-Area-Diff: In this variant, the tracker outputs trajectory points for each frame within the horizon, and SAM predicts masks based on these trajectories.
+Reinitialization happens at the frame within the horizon that has the mean mask area among the non-empty masks predicted by SAM.
+  3. (C)
+    Reinit-on-Similar-Mask-Area: This method triggers reinitialization when the mask area is similar to the initial mask area.
+  4. (D)
+    Reinit-on-Similar-Mask-Area-and-Sync-Masks: This variant reinitializes when the mask area for all masks in the batch is similar to the initial mask areas, synchronizing the masks to be tracked from the same timestep.
+This synchronization allows for the use of negative points from other masks when querying SAM.
+From our ablation investigations, we found the (A) Reinit-on-Horizon-and-Sync-Masks strategy to be effective with PIPS as the point tracker and (B) Reinit-at-Median-of-Area-Diff with CoTracker.
+Besides the point tracker used, the choice of reinitialization method may depend on the specific validation subset and the degree of hyperparameter tuning involved.
+Note that we always use reinitialization along with negative points.
+Appendix H Additional Qualitative Results
+We show additional visualizations on DAVIS videos in Fig. 8. We show failure cases on clips from the anime-influenced series “Avatar: The Last Airbender” in Fig. 9.
+Appendix I VOS Evaluation Details
 When evaluating on VOS, we use the provided ground truth mask for the first frame to sample the query points required by our method.
@@
 For all datasets, we use the full-resolution data and resize it to the longest side of $1024$ to match SAM’s input resolution.
-VIS Evaluation.
+Appendix J VIS Evaluation Details
 For evaluating our method on the VIS task, we leverage SAM’s automatic mask generation capacity to generate up to 100 mask proposals for the initial frame.
-We then propagate these proposed masks throughout the entire video sequence using our method.
-We evaluate TAM [40], a concurrent method we compare against, in the same manner.
-Our mask proposal generation process is currently simplistic and does not create any proposals for subsequent video frames.
-Consequently, it cannot identify objects that emerge in later frames, placing it at a disadvantage compared to VIS methods that are capable of doing so.
-Despite this limitation, our approach provides a consistent platform for comparing zero-shot methods in terms of how effectively they propagate diverse mask proposals from the first frame.
-4.3 Ablation Study
-We conducted detailed ablation experiments on the DAVIS $2017$ validation subset to validate various components and designs of SAM-PT.
-We employed SAM’s ViT-H as the backbone, for all tests.
-Each aspect was examined sequentially, integrating the optimal settings obtained from prior experiments.
-To ensure statistical soundness, multiple iterations of each experiment were carried out (between 4 and 12 runs per setup), with findings represented as mean and standard deviation across these runs.
-While these results provide insight, there may be a risk of overfitting due to our limited validation dataset.
-While we endeavored to maintain a consistent evaluation protocol, future research should aim for a larger validation set, possibly derived from the YouTube VOS $2018$ train dataset, to mitigate this concern.
-Figure 5: Query Point Sampling Ablation Study.
-We report mean $\mathcal{J\&F}$ scores along with the standard deviation on the validation subset of DAVIS $2017$ for different number of positive points per mask and different point selection methods.
-Using $8$ positive points per mask leads to a $40$-point performance boost compared to using a single point.
-Given $8$ positive points, K-Medoids and Shi-Tomasi perform comparably well as point selection methods.
-Query Point Sampling
-Fig.
-5 illustrates that the number of positive points and the choice of point selection methods significantly influence performance.
-Using 8 points per mask showed a remarkable 40-point performance enhancement compared to a single point.
-This substantiates the argument that a single positive point is inadequate for prompting SAM as it often results in the segmentation of partial objects only.
-Among the point selection methods, K-Medoids and Shi-Tomasi produced comparable results, with a slight preference towards K-Medoids owing to its marginally higher mean score and resilience to the number of positive points per mask.
-Table 2: Point Tracker and SAM Configuration Ablation Study Results.
-Using the best parameters from the ablation study in Fig.
-5, we report the mean performance (with standard deviation) on the validation subset of DAVIS $2017$ to study the impact of (a) different point trackers, (b) the number of negative points per mask, (c) the use of iterative refinement and (d) patch similarity filtering.
-We find that the best configuration uses PIPS as the point tracker, $1$ negative point per mask, $12$ refinement iterations, and no patch similarity filtering.
-PT: point tracker.
-NP: negative points per mask.
-IRI: iterative refinement iterations.
-PS: point similarity filtering threshold.
-SAM-PT Configuration DAVIS $\mathbf{2017}$ Validation [27]
-[…]
-(a) point tracker
-RAFT [30] 00 00 $63.0\pm 0.6$ $60.7\pm 0.6$ $65.4\pm 0.5$
-SuperGlue [28] 00 00 $21.7\pm 2.8$ $19.6\pm 2.1$ $23.8\pm 3.4$
-SuperGlue [28] 00 $3$ $28.4\pm 3.1$ $24.7\pm 2.4$ $32.0\pm 3.8$
-TapNet [9] 00 00 $60.9\pm 0.2$ $58.2\pm 0.3$ $63.5\pm 0.2$
-PIPS [11] 00 00 $\mathbf{72.3}\pm 1.2$ $\mathbf{70.4}\pm 1.3$ $\mathbf{74.3}\pm 1.1$ $\mathbf{+9.3}$
-(b) negative points per mask
-PIPS 00 00 $72.3\pm 1.2$ $70.4\pm 1.3$ $74.3\pm 1.1$
-PIPS $1$ 00 $\mathbf{74.1}\pm 0.7$ $\mathbf{72.1}\pm 0.6$ $\mathbf{76.1}\pm 0.7$ $\mathbf{+1.8}$
-PIPS $8$ 00 $74.0\pm 0.8$ $71.9\pm 0.8$ $76.0\pm 0.9$
-PIPS $16$ 00 $73.4\pm 0.6$ $71.4\pm 0.6$ $75.3\pm 0.6$
-PIPS $72$ 00 $72.2\pm 0.4$ $70.3\pm 0.4$ $74.0\pm 0.4$
-(c) iterative refinement iterations
-PIPS $1$ 00 $74.1\pm 0.7$ $72.1\pm 0.6$ $76.1\pm 0.7$
-PIPS $1$ $1$ $75.7\pm 0.7$ $73.4\pm 0.7$ $78.1\pm 0.6$
-PIPS $1$ $3$ $76.0\pm 0.6$ $73.4\pm 0.7$ $78.6\pm 0.7$
-PIPS $1$ $12$ $\mathbf{76.3}\pm 0.6$ $\mathbf{73.6}\pm 0.6$ $\mathbf{78.9}\pm 0.6$ $\mathbf{+2.2}$
-(d) patch similarity filtering
-PIPS $1$ $12$ $\mathbf{76.3}\pm 0.6$ $\mathbf{73.6}\pm 0.6$ $\mathbf{78.9}\pm 0.6$ none
-PIPS $1$ $12$ $0.002$ $72.7\pm 2.0$ $70.2\pm 1.8$ $75.2\pm 2.1$
-PIPS $1$ $12$ $0.01$ $70.7\pm 2.0$ $68.3\pm 1.8$ $73.2\pm 2.1$
-Point Tracking.
-Tab.
-2a shows that PIPS [11] demonstrated superior performance over TapNet [9], SuperGlue [28], and RAFT [30].
-TapNet’s limitations stem from its lack of effective time consistency and its training on 256x256 images, which hampered its performance with higher-resolution images.
-SuperGlue, while proficient in matching sparse features across rigid scenes, grapples with effectively matching points from the reference frame in dynamic scenes, particularly under object deformations.
-RAFT, being an optical flow model, faced difficulties handling occlusions.
-Although PIPS’s prior use in our experiments may have offered some hyperparameter advantages, its superior performance is primarily attributable to its more robust design that emphasizes trajectory modeling over eight subsequent frames.
-This approach fosters the generation of coherent point trajectories and enhances occlusion detection.
-Negative Points.
-Tab.
-2b highlights that incorporating negative points had a favorable impact, particularly in reducing segmentation errors when points deviated from the target object.
-The addition of negative points empowered SAM to better handle the point trackers’ failure cases, leading to improved segmentation and a 1.8-point enhancement over the non-use of negative points.
-Note that throughout all experiments, we always used the mixed point sampling method for sampling negative points which amounts to using random sampling when there is only one negative point per mask.
-Iterative Refinement.
-The iterative refinement approach contributed to higher-quality masks and mitigated the impact of artifacts in SAM’s output.
-Tab.
-2c displays that this yielded an improvement of 2.2 points over the non-refinement approach.
-Patch Similarity.
-Our initial findings in Tab.
-2d suggest that using patch similarity to filter unreliable tracking points was overly restrictive in our context, leading to substantial deletion of points.
-Although it did not prove beneficial in our current setup, this aspect certainly warrants further exploration, particularly in scenarios involving point re-initialization.
-Reinitialization.
-Fig.
-6 presents the performance of different reinitialization variants.
-In Tab.
-5 and Tab.
-6, we also show it brings 2.5 and 2.0 points improvements on MOSE and UVO benchmarks respectively.
-The re-initialization process enhanced robustness against points falling off objects.
-By reinitializing all points based on the current mask prediction, we account for errors in point tracker outputs by discarding incorrect points and starting fresh from the current mask prediction.
-However, this assumes that we trust the currently outputted mask, which may not always be the case and sometimes leads to failures.
-Figure 6: Reinitialization Ablation Study.
-Mean $\mathcal{J\&F}$ scores (with std.
-dev.) on the validation subset of DAVIS $2017$ for different reinitialization variants and configurations.
-The best result is achieved by reinitialization variant $A$, 12 refinement iterations, and 72 negative points per mask.
-The reinitialization variants differ in how the timestep at which the points get reinitialized is chosen, see Appendix A for more details.
-Although using reinitialization improves the performance only marginally on the validation subset, this strategy demonstrates substantial improvement on the MOSE $2023$ (Tab.
-5) and UVO (Tab.
-6) datasets.
-In summary, our best-performing SAM-PT model employs K-Medoids for point selection with 8 points per mask, PIPS for point tracking, a single negative point per mask, and employs 12 iterations for iterative refinement without patch similarity filtering.
-Meanwhile, using reinitialization achieved optimum performance with 12 refinement iterations and 72 negative points per mask.
-4.4 Comparison with State-of-the-art Methods
-All reported results were computed with official tools or official evaluation servers.
-Sec.
-4.4.1 reports Video Object Segmentation results, including qualitative results on unseen web videos.
-Sec.
-4.4.2 reports Video Instance Segmentation results.
-4.4.1 Video Object Segmentation
-Table 3: Quantitative results in semi-supervised VOS on the validation subset of DAVIS $2017$.
-DAVIS $\mathbf{2017}$ Validation [27]
-[…]
-(a) trained on video segmentation data
-AGSS [20] $67.4$ $64.9$ $69.9$
-AGAME [13] $70.0$ $67.2$ $72.7$
-AFB-URR [18] $74.6$ $73.0$ $76.1$
-STM [25] $81.8$ $79.2$ $84.3$
-SWEM [21] $84.3$ $81.2$ $87.4$
-RDE [17] $86.1$ $82.1$ $90.0$
-SwinB-DeAOT-L [41] $86.2$ $83.1$ $89.2$
-XMem [4] $87.7$ $84.0$ $91.4$
-(b) not trained on video segmentation data (zero-shot)
-Painter [35] $34.6$ $28.5$ $40.8$
-DINO [2] $71.4$ $67.9$ $74.9$
-SegGPT [36] 75.675.6{75.6}75.6 72.572.5{72.5}72.5 78.678.6{78.6}78.6
-PerSAM-F [45] $71.9$ $69.0$ $74.8$
-SAM-PT (ours) $76.3\pm 0.6$ $73.6\pm 0.6$ ${78.9}\pm 0.6$
-SAM-PT-reinit (ours) ${76.6}\pm 0.7$ ${74.4}\pm 0.8$ ${78.9}\pm 0.6$
-HQ-SAM-PT [15] (ours) $\mathbf{77.2}\pm 0.5$ $74.7\pm 0.5$ $\mathbf{79.8}\pm 0.4$
-HQ-SAM-PT-reinit [15] (ours) $77.0\pm 0.7$ $\mathbf{74.8}\pm 0.8$ $79.2\pm 0.6$
-Performance Overview.
-Our proposed method outperforms others that have not been trained on any video object segmentation data on the DAVIS $2017$ dataset, as reflected in Tab.
-3.
-A mean $\mathcal{J\&F}$ score of $76.6$ points exceeds the PerSAM-F by $4.7$ points and the SegGPT generalist model by a single point.
-The experiments were repeated 8 times for statistical robustness, and we report the mean and standard deviation of our method’s performance.
-We also outperform PerSAM-F on the YouTube-VOS $2018$ and MOSE $2023$ datasets, achieving mean scores of $67.0$ and $41.0$ as shown in Tabs.
-4 and 5.
-However, with different mask training data, our performance falls short when compared to SegGPT on the two datasets.
-(a) Good cases for SAM-PT
-(b) Good cases for SAM-PT-reinit
-(c) Failure cases for SAM-PT
-Figure 7: Visualization of SAM-PT on the DAVIS 2017 Validation [27] dataset.
-Given the first-frame masks, we sample 8 positive points and either 1 (SAM-PT) or 72 (SAM-PT-reinit) negative points per object to initialize SAM-PT and SAM-PT-reinit.
-Circles denote positive points, crosses denote negative points, and red symbols (circle and cross) denote that the point was predicted to be occluded.
-Figure 8: Successful segmentation using SAM-PT on short clips from “Avatar: The Last Airbender”.
-Although our method has never seen data from Avatar, an anime-influenced animated television series, it segments and tracks various objects in short clips.
+We use the same initial masks to assess TAM [46] for a fair comparison.
+Our current approach does not generate new proposals in subsequent frames, thus it cannot detect new objects appearing after the first frame, unlike fully-fledged VIS methods.
+However, this setup allows for a straightforward comparison of zero-shot capabilities in mask propagation from the initial frame.
+Appendix K BDD100K VOS Dataset Creation
+BDD100K is a large open driving video dataset with 100K videos and 10 tasks to evaluate the progress of image recognition algorithms in autonomous driving.
+It includes a variety of geographic, environmental, and weather conditions.
+We convert its annotations from the Multi-Object Tracking and Segmentation (MOTS) section into semi-supervised VOS annotations.
+This section has $154$, $32$, and $37$ videos for train, validation, and test sets, totaling $25$K instances and $480$K masks.
+To convert the annotations, we take the ground truth mask of the first appearance of each object in the videos.
+This mask will be given as input to semi-supervised VOS methods which then need to predict the masks for the remaining video frames.
+In converting the validation subset of $32$ videos and $4566$ object tracks, we create $61$ semi-supervised VOS datapoints of up to $100$ objects per video.
+We limit the number of objects per video for implementation simplicity since most VOS methods expect a small number of objects per video.
+For example, in the DAVIS validation subset, the maximum number of objects per video is 5.
+During the conversion, we additionally remove the instances marked as “ignored” or “crowd” in the MOTS annotations.
+Appendix L Interactive Point-Based Video Segmentation Details
+Interactive point-based video segmentation aims to refine segmentation masks with minimal user input while optimizing the Intersection over Union (IoU) with the ground truth.
+To evaluate the responsiveness of SAM-PT to simulated human input, we benchmark against a SAM-only baseline (Algorithm 1) and compare it with both an online (Algorithm 2) and offline (Algorithm 3) interactive SAM-PT method:
+  1. 1.
+    Non-tracking method (SAM only): As a baseline, this method mimics user interactions by selecting points on each frame without any point tracking, effectively emulating the process of manual, frame-by-frame annotation using the standalone SAM model.
+  2. 2.
+    Online method: This approach models a user going through the video sequentially a single time, making corrective per-frame interactions to achieve an IoU of at least 95% with the ground truth mask for each frame.
+These corrections include the addition or removal of points and are propagated to subsequent frames via point tracking.
+The user may opt to skip frames that already appear to be well-annotated or that cannot be annotated sufficiently well within the available interaction budget.
+  3. 3.
+    Offline method (Checkpoint Strategy): This method takes a multi-pass approach, incrementally aiming for higher IoU thresholds with each pass through the video.
+Checkpoints are saved after each pass, and the latest checkpoint given the interaction budget is returned, allowing for a more global optimization approach.
+Interactions are defined as the act of adding or removing a point and are executed as described in Algorithm 4.
+To maintain simplicity in our evaluation, skipping frames while progressing through a video is not counted as an interaction.
+Algorithm 1 Non-tracking (SAM only) Method.
+1:   Input: rgbs, gt_masks, int_per_frame
+2:   Output: pred_masks
+3:   point_memory $\leftarrow$ emptyPointMemory()
+4:   for i in frames do
+5:      pmf $\leftarrow$ emptyFramePointMemory()
+6:      for j in int_per_frame do
+7:         m $\leftarrow$ predMask(rgbs[i], pmf)
+8:         performInt(i, m, gt_masks[i], pmf)
+9:      end for
+10:      addToPointMemory(point_memory, i, pmf)
+11:   end for
+12:   return  predAllMasks(rgbs, point_memory)
+Algorithm 2 Online Method.
+1:   Input: rgbs, gt_masks
+2:   Output: pred_masks
+3:   max_int $\leftarrow 300$
+4:   max_int_per_frame $\leftarrow 3$
+5:   threshold $\leftarrow 0.95$
+6:   point_memory $\leftarrow$ selectFirstPoint(gt_masks[0])
+7:   max_int $\leftarrow$ max_int - 1
+8:   for i in frames do
+9:      m $\leftarrow$ predMask(rgbs[i], point_memory)
+10:      IoU $\leftarrow$ calculateIoU(m, gt_masks[i])
+11:      if IoU $\geq$ threshold then
+12:         continue
+13:      end if
+14:      performInt(i, m, gt_masks[i], point_memory)
+15:      max_int $\leftarrow$ max_int - 1
+16:      if max_int $\leq 0$ then
+17:         break
+18:      end if
+19:   end for
+20:   return  predAllMasks(rgbs, point_memory)
+Algorithm 3 Offline Method (Checkpoint Strategy).
+1:   Input: rgbs, gt_masks
+2:   Output: pred_masks
+3:   max_int $\leftarrow 300$
+4:   max_int_per_frame $\leftarrow 3$
+5:   int_iou_thresholds $\leftarrow[0.10,0.20,\ldots,0.95]$
+6:   point_memory $\leftarrow$ selectFirstPoint(gt_masks[0])
+7:   max_int $\leftarrow$ max_int - 1
+8:   best_ckpt $\leftarrow$ predAllMasks(rgbs, point_memory)
+9:   for threshold in int_iou_thresholds do
+10:      for i in frames do
+11:         for j in max_int_per_frame do
+12:            m $\leftarrow$ predMask(rgbs[i], point_memory)
+13:            IoU $\leftarrow$ calculateIoU(m, gt_masks[i])
+14:            if IoU $\geq$ threshold then
+15:               break
+16:            end if
+17:            performInt(i, m, gt_masks[i], point_memory)
+18:            max_int $\leftarrow$ max_int - 1
+19:         end for
+20:      end for
+21:      if max_int $\geq 0$ then
+22:         best_ckpt $\leftarrow$ predAllMasks(rgbs, point_memory)
+23:      else
+24:         break
+25:      end if
+26:   end for
+27:   return  best_ckpt
+Algorithm 4 Perform Interaction.
+1:   Input: frame_idx, m, gt_m, point_memory
+2:   i $\leftarrow$ frame_idx
+3:   tp_mask $\leftarrow$ m $\land$ gt_m
+4:   tn_mask $\leftarrow$ $\lnot$¬m $\land$ $\lnot$¬gt_m
+5:   fp_mask $\leftarrow$ m $\land$ $\lnot$¬gt_m
+6:   fn_mask $\leftarrow$ $\lnot$¬m $\land$ gt_m
+7:   pos_points $\leftarrow$ getPosPoints(point_memory, i)
+8:   neg_points $\leftarrow$ getNegPoints(point_memory, i)
+9:   if any neg_points in fn_mask then
+10:      p $\leftarrow$ firstIncorrectNegPoint(fn_mask, neg_points)
+11:      removePointAndItsFuture(point_memory, i, p)
+12:   else if any pos_points in fp_mask then
+13:      p $\leftarrow$ firstIncorrectPosPoint(fp_mask, pos_points)
+14:      removePointAndItsFuture(point_memory, i, p)
+15:   else
+16:      if sum(fn_mask) $>$ sum(fp_mask) then
+17:         [x, y] $\leftarrow$ extractPoint(fn_mask)
+18:         lbl $\leftarrow$ positive
+19:      else
+20:         [x, y] $\leftarrow$ extractPoint(fp_mask)
+21:         lbl $\leftarrow$ negative
+22:      end if
+23:      addToFrameAndFuture(point_memory, i, lbl, x, y)
+24:   end if
+(a) Successful cases for SAM-PT.
+(b) Failure cases for SAM-PT.
+Figure 8: Additional visualization of SAM-PT on videos from the DAVIS 2017 [30] validation subset, including (a) successful cases and (b) failure cases.
+Circles denote positive points and crosses denote negative points.
+Red symbols indicate occlusion prediction.
 Figure 9: Challenging scenarios for SAM-PT on short clips from “Avatar: The Last Airbender”.
 These cases illustrate instances where our model struggles when faced with point tracking failures that are the result of incorrectly predicting the point at a similar-looking segment or when faced with object occlusions and disappearing objects.
-Qualitative Analysis.
-Visualizations of successful video segmentation on DAVIS $2017$ for SAM-PT and SAM-PT-reinit can be seen in Fig.
-6(a) and Fig.
-6(b) respectively.
-Notably, Fig.
-8 presents successful video segmentation on unseen web videos – clips from the “Avatar: The Last Airbender” anime-influenced animated television series, demonstrating the zero-shot capabilities of our method.
-Limitations and Challenges.
-Despite the competitive zero-shot performance, certain limitations persist, primarily due to the limitations of our point tracker in handling occlusion, small objects, motion blur, and re-identification.
-In such scenarios, the point tracker’s errors propagate into future video frames.
-Fig.
-6(c) illustrates these problematic instances on DAVIS $2017$, while Fig.
-9 presents additional cases on “Avatar: The Last Airbender” clips.
-Although using point re-initialization and negative points somewhat alleviates the failures of the point tracker, they still prevent the performance from being on par with methods trained on video data.
-Table 4: Quantitative results in semi-supervised VOS on the validation subset of YouTube-VOS $2018$.
-Metrics are reported separately for “seen” and “unseen” classes, with $\mathcal{G}$ being the overall average score over the metrics.
-Note that SegGPT and SAM-PT adopt completely different training data.
-♠: our reproduced result using the official code of [45].
-YouTube-VOS $\mathbf{2018}$ Validation [38]
-Method $\mathcal{G}$ $\mathcal{J}_{s}$ $\mathcal{F}_{s}$ $\mathcal{J}_{u}$ $\mathcal{F}_{u}$
-(a) trained on video segmentation data
-AGAME [13] $66.0$ $66.9$ $61.2$
-AGSS [20] $71.3$ $71.3$ $65.5$ $75.2$ $73.1$
-STM [25] $79.4$ $79.7$ $84.2$ $72.8$ $80.9$
-AFB-URR [18] $79.6$ $78.8$ $83.1$ $74.1$ $82.6$
-RDE [17] $83.3$ $81.9$ $86.3$ $78.0$ $86.9$
-SWEM [21] $82.8$ $82.4$ $86.9$ $77.1$ $85.0$
-XMem [4] $86.1$ $85.1$ $89.8$ $80.3$ $89.2$
-SwinB-DeAOT-L [41] $86.2$ $85.6$ $90.6$ $80.0$ $88.4$
-(b) not trained on video segmentation data (zero-shot)
-Painter [35] $24.1$ $27.6$ $35.8$ $14.3$ $18.7$
-SegGPT [36] $\mathbf{74.7}$ $\mathbf{75.1}$ $\mathbf{80.2}$ $\mathbf{67.4}$ $\mathbf{75.9}$
-PerSAM-F ♠ [45] $54.4$ $53.9$ $56.4$ $50.7$ $56.6$
-SAM-PT (ours) $67.0\pm 0.3$ $68.6\pm 0.2$ $71.2\pm 0.1$ $61.0\pm 0.5$ $67.4\pm 0.4$
-SAM-PT-reinit (ours) ${67.5}\pm 0.2$ ${69.0}\pm 0.4$ ${69.9}\pm 0.3$ ${63.2}\pm 0.4$ ${67.8}\pm 0.5$
-Table 5: Quantitative results in semi-supervised VOS on the validation subset of MOSE $2023$ [8]. ♠: our reproduced result using the official code of [45].
-MOSE $\mathbf{2023}$ Validation [8]
-[…]
-(a) trained on video segmentation data
-RDE [17] $48.8$ $44.6$ $52.9$
-SWEM [21] $50.9$ $46.8$ $54.9$
-XMem [4] $57.6$ $53.3$ $62.0$
-DeAOT [41] $59.4$ $55.1$ $63.8$
-(b) not trained on video segmentation data (zero-shot)
-Painter [35] $14.5$ $10.4$ $18.5$
-SegGPT [36] $\mathbf{45.1}$ $\mathbf{42.2}$ $\mathbf{48.0}$
-PerSAM-F ♠ [45] $23.3$ $19.8$ $26.8$
-SAM-PT (ours) $38.5\pm 0.2$ $34.9\pm 0.3$ $42.1\pm 0.2$
-SAM-PT-reinit (ours) ${41.0}\pm 0.5$ ${38.5}\pm 0.5$ ${43.5}\pm 0.5$
-4.4.2 Video Instance Segmentation
-Results and Analysis.
-Given the same mask proposals, SAM-PT outperforms TAM [40] significantly even though SAM-PT was not trained on any video segmentation data.
-TAM is a concurrent approach combining SAM and XMem [4], where XMem was pre-trained on BL30K [5] and trained on DAVIS and YouTube-VOS, but not on UVO.
-On the other hand, SAM-PT combines SAM with the PIPS point tracking method, both of which have not been trained on video segmentation tasks.
-5 Conclusion
-We present SAM-PT, an innovative solution that extends SAM’s segmentation ability from static images to dynamic videos.
-Integrated with long-term point trackers, our approach demonstrates strong performance across several benchmarks including DAVIS, YouTube-VOS, MOSE, and UVO.
-While our method has limitations such as difficulty handling occlusions, small objects, and motion blur, and inconsistencies in mask predictions, it contributes a simple and effective new point-based perspective to video object segmentation research.
-By illustrating a promising way to extend foundational models like SAM into the video domain, our research provides a potential pathway for advancements in diverse applications from autonomous driving to video labeling.
-Furthermore, the future incorporation of more advanced point trackers can enhance the performance of SAM-PT.
-[…]
-Method Propagation $\mathbf{AR100}$ $\mathbf{ARs}$ $\mathbf{ARm}$ $\mathbf{ARl}$ $\mathbf{AP}$
-(a) trained on video segmentation data, including UVO’s training subset
-Mask2Former VIS [44] N/A $35.4$- -- -- $27.3$
-ROVIS [44] N/A $41.2$- -- -- $32.7$
-(b) trained on video segmentation data
-TAM [40] Feature Matching $24.1$ $21.1$ $32.9$ $31.1$ $1.7$
-(c) not trained on video segmentation data (zero-shot)
-SAM-PT (ours) Points Prompting $28.8\pm 0.1$ $23.3\pm 0.1$ $40.8\pm 0.2$ $48.3\pm 0.6$ $\mathbf{6.7}\pm 0.2$
-SAM-PT-reinit (ours) Points Prompting $\mathbf{30.8}$ $\mathbf{25.1}$ $\mathbf{44.1}$ $\mathbf{49.2}$ $6.5$
-Appendix A Point Tracking Reinitialization
-In our SAM-PT-reinit method, we introduce a reinitialization strategy.
-Here, the point tracker begins anew after every $h$ frames, where $h$ represents a pre-set tracking horizon (e.g., 8 frames), or is dynamically determined based on SAM’s mask predictions for each timestep within the horizon (e.g., using most-similar-mask-area heuristics).
-Upon reaching this horizon, the query points given to the tracker are reinitialized according to the mask prediction SAM outputted at the horizon frame.
-While this method may increase the computational load (especially if some of SAM’s computed masks are disregarded), it demonstrates substantial performance improvement in demanding video sequences, such as those in the MOSE dataset.
-We explored four reinitialization methods, each varying in how they compute the value of $h$:
-  1. (A)
-    Reinit-on-Horizon-and-Sync-Masks: This straightforward variant reinitializes points after a fixed number of frames (e.g., every 8 frames).
-However, it may stumble if the mask is absent at the reinitialization timestep.
-Despite this potential pitfall, it operates at the same speed as methods that do not employ reinitialization.
-  2. (B)
-    Reinit-at-Median-of-Area-Diff: In this variant, the tracker outputs trajectory points for each frame within the horizon, and SAM predicts masks based on these trajectories.
-Reinitialization happens at the frame within the horizon that has the mean mask area among the non-empty masks predicted by SAM.
-Notably, this approach may be significantly slower than methods without reinitialization, as it may reject several SAM masks (e.g., out of 8 computed masks, reinitialization might occur on the second one, necessitating recomputation of the remaining 6 masks in the next step).
-  3. (C)
-    Reinit-on-Similar-Mask-Area: This method triggers reinitialization when the mask area is similar to the initial mask area, causing it to be several times slower than methods without reinitialization.
-  4. (D)
-    Reinit-on-Similar-Mask-Area-and-Sync-Masks: This variant reinitializes when the mask area for all masks in the batch is similar to the initial mask areas, synchronizing the masks to be tracked from the same timestep.
-This synchronization allows for the use of negative points from other masks when querying SAM, but it also runs several times slower than methods without reinitialization.
-From our investigations, we found the (A) Reinit-on-Horizon-and-Sync-Masks strategy to be the most effective, as indicated by its superior performance on the DAVIS $2017$ validation subset.
-The choice of reinitialization method may depend on the specific validation subset and the degree of hyperparameter tuning involved.
-Note that we have always used reinitialization along with negative points.
-A.1 Computational Cost and Speed Optimization
-The introduction of reinitialization in SAM-PT-reinit comes with a trade-off: it slows down the inference speed by a factor of 2 to 8, depending on the reinitialization method and parameters used.
-The major bottleneck is the invocation of SAM’s backbone for each video frame.
-We propose caching the backbone outputs for unprocessed video frames as a possible solution to mitigate this slowdown.
-This strategy requires storing embeddings for all video frames in the working memory but offers the potential for significant speedup, particularly useful for applications requiring faster inference.
-Appendix B More DAVIS Subsets
-We report results on DAVIS $2016$ Validation and DAVIS $2017$ Test-dev in Tab. 7.
-Table 7: Quantitative results in semi-supervised video object segmentation on additional subsets of DAVIS. ♠: our reproduced result using the official code of [45].
-DAVIS $\mathbf{2016}$ Validation [27] DAVIS $\mathbf{2017}$ Test-dev [27]
-[…]
-PerSAM-F ♠ [45] $74.8$ $74.5$ $75.0$ $47.6$ $45.5$ $49.7$
-SegGPT [36] $82.3$ $81.8$ $82.8$- - -
-SAM-PT $\mathbf{83.1}\pm 1.5$ $\mathbf{83.0}\pm 0.8$ $\mathbf{83.0}\pm 1.1$ $\mathbf{62.7}\pm 0.5$ $\mathbf{59.4}\pm 0.6$ $\mathbf{66.1}\pm 0.4$
-SAM-PT-reinit $80.2\pm 0.6$ $80.3\pm 0.6$ $80.0\pm 0.6$ $61.5\pm 1.1$ $59.3\pm 1.0$ $63.8\pm 1.2$
-Appendix C Per-sequence DAVIS 2017 Validation Results
-See figure Fig. 10 for per-sequence DAVIS 2017 Validation results. For exact numbers and tables, check our GitHub experiment summaries or the Wandb project.
-(a) number of positive points per mask
-(b) point selection methods
-(c) point tracking module
-(d) number of negative points per mask
-(e) number of iterative refinement iterations
-(f) patch similarity filtering threshold
-Figure 10: Per-sequence Ablation Results.
-The charts plot the median J-Mean result of different ablation result experiments along with $95\%$ bootstrapped confidence intervals, based on 1000 bootstrap samples.